### v1

# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

Abstract
We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data. vec2vec finds a near-perfect alignment, but it is expensive and unstable. We present mini-vec2vec, a simple and efficient alternative that requires substantially lower computational cost and is highly robust. Moreover, the learned mapping is a linear transformation. Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement. Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results. The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
Algorithm 1 Unsupervised Embedding Alignment
1:Embedding matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$, $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$
2:Number of clusters $c$, neighbors $k$, iterations $T$, smoothing parameter $\alpha$, runs $s$, clusters $c^{\prime}$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Linear transformation matrix $\mathbf{W}$
4:Normalize embeddings: $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}\leftarrow\text{center and normalize to unit sphere}$
5:Create pseudo-pairs: $\mathcal{C}\leftarrow\operatorname{AnchorAlignment}(\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B},c,k,s)$ $\triangleright$ See Alg. 2
6:Estimate initial mapping: $\mathbf{W}\leftarrow\text{Procrustes analysis on }\mathcal{C}$
7:Refine-1: Apply matching-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 3
8:Refine-2: Apply clustering-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 4
9:return $\mathbf{W}$
1 Introduction
Representation learning has revolutionized natural language processing by producing embedding models that capture rich semantic structure in high-dimensional vector spaces. These learned representations encode complex linguistic relationships, enabling downstream tasks such as semantic similarity and information retrieval. However, different models, training procedures, or even different random initializations of the same architecture can produce embedding spaces that, while semantically meaningful, are not directly comparable due to arbitrary rotations, reflections, and translations in their coordinate systems.
The problem of aligning embedding spaces without parallel supervision – that is, without access to pairs of embeddings that represent identical semantic content – has emerged as a fundamental challenge with far-reaching implications. Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (Conneau et al., 2018; Zhang et al., 2017), cross-modal alignment between text and image representations (Maniparambil et al., 2024), and privacy concerns (Song and Raghunathan, 2020; Morris et al., 2023).
The recent vec2vec method (Jha et al., 2025) demonstrated that unsupervised alignment is indeed possible using CycleGAN (Zhu et al., 2020), an adversarial training framework. Their approach is theoretically grounded in the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces. Under this hypothesis, semantic relationships are preserved across models in terms of relative distances and angles, suggesting that alignment should be achievable through geometric transformations.
While the vec2vec approach has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability. GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017). These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
Our point of departure is still universal geometry, but we pursue a different approach. We draw inspiration from the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations. Unfortunately, stitching requires access to pairs of matching representations. Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training. Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision. To this end, we utilize the Relative Representation framework of Moschella et al. (2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
Our contribution is mini-vec2vec, a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (running on commodity CPU hardware), training stability (deterministic components with controllable stochastic elements), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples). Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
2 Preliminaries
The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales. Early work by Olah (2015) revealed striking structural regularities in representations of different convolutional networks, showing distances between object representations are roughly the same regardless of architecture. These findings suggested that the geometric organization of learned features might be more universal than previously assumed. The Platonic Representation Hypothesis, formally articulated by Huh et al. (2024), represents the most comprehensive theoretical framework for understanding universal geometry. The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain.
Moschella et al. (2023), inspired by Olah (2015), utilized these observations by proposing relative representations as a mathematically principled framework for comparing embedding spaces. Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable. This work provided both theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
Given a set of anchor points $\{\mathbf{a}_{i}\}_{i=1}^{k}$ in an embedding space, the relative representation of any point $\mathbf{x}$ is defined as:
$\mathbf{r}(\mathbf{x})=[f(\mathbf{x},\mathbf{a}_{1}),f(\mathbf{x},\mathbf{a}_{2}),\ldots,f(\mathbf{x},\mathbf{a}_{k})]^{T},$
where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity). Two representations from different models will have close relative representations if they represent similar objects.
Algorithm 2 Algorithm for creating pseudo-pairs using noisy anchors
1:Normalized embeddings $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, clusters $c$, neighbors $k$, runs $s$
2:Set of pseudo-parallel pairs $\mathcal{C}$
3:Initialize collections: $\mathcal{R}_{A}\leftarrow[],\mathcal{R}_{B}\leftarrow[]$
4:for $i=1$ to $s$ do $\triangleright$ Multiple runs for robustness
5:  Cluster both spaces: $\{\mathbf{c}^{A}_{i,j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{A})$, $\{\mathbf{c}^{B}_{i,j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{B})$
6:  Compute similarities: $\mathbf{S}^{A},\mathbf{S}^{B}\leftarrow\text{cosine similarities between centroids}$
7:  Find correspondence: $\mathbf{P}^{*}_{i}\leftarrow\text{solve QAP}(\mathbf{S}^{A},\mathbf{S}^{B})$ using multiple 2-OPT runs
8:  Build relative representations: $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}\leftarrow\text{cosine similarities to anchors}$
9:  Store: append $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}$ to $\mathcal{R}_{A},\mathcal{R}_{B}$
10:end for
11:Concatenate: $\mathbf{r}^{A}\leftarrow\text{concatenate all runs in }\mathcal{R}_{A}$
12:Concatenate: $\mathbf{r}^{B}\leftarrow\text{concatenate all runs in }\mathcal{R}_{B}$
[…]
14:Create pseudo-pairs: $\mathcal{C}\leftarrow\left\{(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{j}\mathbf{x}^{B}_{j}):\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})\right\}$
15:return $\mathcal{C}$
3 Problem Formulation
We follow the setup laid out in vec2vec. We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples. Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points. This approach offers key advantages:
  • Generalization: New embeddings can be translated without re-applying the algorithm
  • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
  • Out-of-distribution capability: The parameterized nature enables generalization beyond the training data
The universal geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries where cosine similarities remain approximately invariant across embedding spaces. These geometric similarities enable the identification of this alignment.
Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$. These embeddings represent semantic content from potentially different models, architectures, or training procedures, but we assume they encode similar semantic information from the same underlying distribution $\mathcal{D}$. Unlike more traditional setups, there does not exist a pairing between data points, i.e., there is no overlap between the sentences themselves.
We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector. The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation. Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content. [^1] More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$. We will not take advantage of this flexibility in the paper.
Algorithm 3 Refine-1: Matching-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Iterations $T$, smoothing parameter $\alpha$, neighbors $k^{\prime}$, sample size $n_{s}$
3:Refined transformation matrix $\mathbf{W}$ after matching-based refinement
4:for $t=1$ to $T$ do
5:  Sample embeddings: $\mathbf{X}_{\text{sample}}\leftarrow\text{random sample of }n_{s}\text{ rows from }\hat{\mathbf{X}}_{A}$
6:  Transform samples: $\mathbf{X}_{\text{transformed}}\leftarrow\mathbf{X}_{\text{sample}}\mathbf{W}$
[…]
8:  Average neighbors: $\mathbf{X}_{\text{matched}}\leftarrow\text{mean of neighbors for each sample}$
9:  Construct pseudo-pairs: $\mathcal{C}^{\prime}\leftarrow\{(\mathbf{X}_{\text{sample}},\mathbf{X}_{\text{matched}})\}$
10:  Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
11:  Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
12:end for
Algorithm 4 Refine-2: Clustering-Based Refinement
1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
2:Smoothing parameter $\alpha$, number of clusters $c^{\prime}$
3:Refined transformation matrix $\mathbf{W}$ after clustering-based refinement
4:Cluster space A: $\{\mathbf{c}^{A}_{j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{A},c^{\prime})$
5:Transform centroids: $\{\mathbf{c}^{A\to B}_{j}\}\leftarrow\{\mathbf{c}^{A}_{j}\mathbf{W}\}$
6:Cluster space B with seeds: $\{\mathbf{c}^{B}_{j}\}\leftarrow\text{K-means}(\hat{\mathbf{X}}_{B},c^{\prime},\text{init}=\{\mathbf{c}^{A\to B}_{j}\})$
7:Construct pairs $\mathcal{C}^{\prime}$ by matching cluster centroids
8:Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
9:Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
10:return $\mathbf{W}$
vec2vec mini-vec2vec
$M_{1}$ $M_{2}$ Top-1 $\uparrow$ Rank $\downarrow$ Initial Refine-1 Refine-2 Top-1 $\uparrow$ Rank $\downarrow$
gtr 0.99 1.19 (0.1) 0.25 (0.02) 0.58 (0.00) 0.57 (0.00) 0.99 (0.00) 1.02 (0.00)
gran. stel. 0.98 1.05 (0.0) 0.38 (0.00) 0.59 (0.00) 0.60 (0.00) 0.99 (0.00) 1.03 (0.00)
e5 0.98 1.11 (0.0) 0.28 (0.00) 0.52 (0.00) 0.53 (0.00) 0.99 (0.00) 1.05 (0.00)
gran. 0.99 1.02 (0.0) 0.31 (0.02) 0.58 (0.00) 0.57 (0.00) 1.00 (0.00) 1.01 (0.00)
gtr stel. 0.99 1.03 (0.0) 0.29 (0.00) 0.53 (0.00) 0.54 (0.00) 0.98 (0.00) 1.03 (0.00)
e5 0.84 2.88 (0.2) 0.22 (0.04) 0.48 (0.00) 0.49 (0.00) 0.98 (0.00) 1.06 (0.00)
gran. 0.98 1.08 (0.0) 0.37 (0.02) 0.58 (0.00) 0.59 (0.00) 0.99 (0.00) 1.02 (0.00)
stel. gtr 1.00 1.10 (0.0) 0.28 (0.02) 0.54 (0.00) 0.54 (0.00) 0.96 (0.00) 1.10 (0.00)
e5 1.00 1.00 (0.0) 0.43 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
gran. 0.99 2.20 (0.2) 0.30 (0.00) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.04 (0.00)
e5 gtr 0.82 2.56 (0.0) 0.22 (0.03) 0.49 (0.00) 0.51 (0.00) 0.96 (0.00) 1.10 (0.00)
stel. 1.00 1.00 (0.0) 0.44 (0.01) 0.64 (0.00) 0.65 (0.00) 1.00 (0.00) 1.00 (0.00)
Table 1: Results show mean $\pm$ standard deviation over 3 runs. Bold indicates best performance. Underline indicates nearly equally good performance ($\leq 0.01$ difference). Bold and underline combined indicate better performance, but only within a $0.01$ margin.
4 Methodology
4.1 Motivation
In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$. Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$. This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces. Later on, we will relax this assumption. Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
Our linear approach addresses key limitations of existing methods: GANs are notoriously unstable and hard to train with many hyperparameters, require extensive computational resources and training time, and often need multiple seeds to achieve good solutions. In contrast, we know that in supervised cases (model stitching), there usually exists an affine mapping between latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently. Our method requires significantly less data (60k vs. 1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
Our method resembles Hoshen and Wolf (2018)’s approach for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement. This scheme is very powerful, and we apply a similar method. We find that the approximate matching stage especially requires more delicate treatment in our case, so their approach is not applicable as-is.
The stages incrementally improve alignment quality – each stage builds upon structural correspondences to establish increasingly accurate alignments:
  • Approximate Matching: Find approximate pairs of embeddings from both spaces using landmark points as anchors for relative representations
  • Mapping Estimation: Learn a linear orthogonal mapping from approximate matchings, where linearity and orthogonality provide an inductive bias to smooth out incorrect pairings
  • Iterative Refinement: Use the initial solution as a stepping stone for better matching/mapping, iterating until convergence or budget exhaustion
The overall process is described in Algorithm 1.
4.2 Approximate Matching
Preprocessing.
We first address arbitrary translations and scalings across embedding spaces. Both spaces are centered around their respective means and normalized to the unit hypersphere:
$\boldsymbol{\mu}_{A}$ $=\frac{1}{n_{A}}\mathbf{X}_{A}^{T}\mathbf{1}_{n_{A}},\quad\tilde{\mathbf{X}}_{A}=\mathbf{X}_{A}-\mathbf{1}_{n_{A}}\boldsymbol{\mu}_{A}^{T}$
$\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
Centroid Matching.
To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching. Each run consists of K-means clustering followed by solving a quadratic assignment problem (QAP) for cluster correspondence.
In each run, we first perform K-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$. We then compute structural similarity matrices between cluster centroids:
$S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
We find the optimal matching between cluster centroids by solving the quadratic assignment problem:
$\mathbf{P}^{*}=\operatorname{argmax}_{\mathbf{P}\in\Pi_{k}}\text{Tr}\left(\mathbf{S}_{A}^{T}\mathbf{P}\mathbf{S}_{B}\mathbf{P}^{T}\right)$
where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices. For solving QAP, we use $2$-OPT (running 30 times and choosing the one with the highest alignment score for further robustness).
Relative Representations.
Once we have identified an alignment between centroids, we use the aligned centroids as anchors. For each run within the anchor discovery phase, we represent each embedding through its relationships to anchor points. For embedding $\mathbf{x}^{A}_{i}$ in a given run, we construct:
$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]^{T}$
where $\mathbf{r}^{A}_{i}$ is the relative representation of the $i$-th embedding in space $A$ with respect to the $k$ cluster centroids from that run.
Robustification through Ensembling.
We improve robustness by concatenating relative representations from multiple runs:
$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk}$
This concatenation creates richer structural signatures that are more robust to clustering initialization randomness.
Anchor alignment algorithm is provided in Algorithm 2.
4.3 Mapping Estimation
We construct pseudo-parallel pairs by matching embeddings by similarity in relative space, averaging the $k$ nearest neighbors to reduce noise:
$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\mathbf{x}^{B}_{j}\right):i=1,\ldots,n_{A}\right\}$
where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space. Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$. The orthogonal transformation is obtained via Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
4.4 Iterative Refinement
The initial transformation provides a coarse alignment that we refine with two complementary strategies. This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima. Pseudo-code is provided in Algorithm 3 & 4.
Refine-1: Matching-Based Refinement.
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, and averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.3), with the crucial difference that now the similarity is performed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100) using subsampling to boost computational efficiency.
We use exponential smoothing for transformation updates:
$\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
with $\alpha=0.5$, though the results appear robust to other choices as well.
Refine-2: Clustering-Based Refinement.
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing. [^2] Note that the above justification for Refine-2 is our a priori rationale for using this method, and we do not make any claims as to whether this is the true reason for the ability of this method to improve the alignment. Our only claim is that we observe this step improves the alignment consistently, and despite the fact that the Refine-1 has already converged and no longer shows improvement.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Curiously, two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
5 Experiments
We replicate the in-distribution experiment from Jha et al. (2025). We run the algorithm for all pairs among four sets of text encoders used in Jha et al. (2025): stella, gtr, granite, and e5 (Zhang et al., 2025; Ni et al., 2021; Awasthy et al., 2025; Wang et al., 2024). Experiments use K-Means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT (Croes, 1958).
We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime, and repeated three times for each pair to demonstrate consistency. We compare the results to the numbers reported in Jha et al. (2025).
Metrics.
We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\text{NN}(\mathbf{W}\mathbf{a}_{i})=\mathbf{b}_{i}),$
where $\delta$ is the indicator function, $\text{NN}(\cdot)$ is the nearest neighbor of source embedding $\mathbf{a}_{i}$ in the target space, and $\mathbf{b}_{i}$ is its true corresponding target embedding. The second is the Average Rank, the mean position of the true match in the sorted list of distances from the query to all targets,
$\text{AvgRank}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}(\mathbf{b}_{i}|\mathbf{a}_{i}).$
We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages. [^3] Cosine similarity in itself is indicative, but it can also be misleading. Cosine similarity in mini-vec2vec underestimates the similarity, in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
Hyperparameters.
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. We use $k=50$ neighbors in Mapping Estimation. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$.
Note this is a rather economical choice of hyperparameters, which we would not have suggested to use a priori. It turned out to work well, regardless. The full hyperparameter configuration used in the experiments is also documented in the code linked.
Results.
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. Our advantage is more noticeable in the rank metric. Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already at the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $e5$ and gtrgtr pair.
We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise. Our results are not meant to show superiority, but rather matching performance for a fraction of the cost – except in the cases where vec2vec converges poorly, in which we observe mini-vec2vec’s superior behavior. Computationally, one run is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
Analysis.
We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape may contain well-connected regions or multiple basins with similar minima. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
6 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. Olah (2015) studied the equivalence of neural networks in relative terms. Moschella et al. (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Huh et al. (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of usupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018). Most of the works rely on training a GAN. They often require a small seed of matched pairs in order to work. Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
7 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, provided that structural correspondences can be estimated accurately.
The success of our approach provides strong empirical evidence for the universal geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).

### v2

@@
 vec2vec finds a near-perfect alignment, but it is expensive and unstable.
-We present mini-vec2vec, a simple and efficient alternative that requires substantially lower computational cost and is highly robust.
+We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost.
 Moreover, the learned mapping is a linear transformation.
@@
 Under this hypothesis, semantic relationships are preserved across models in terms of relative distances and angles, suggesting that alignment should be achievable through geometric transformations.
-While the vec2vec approach has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability.
+While vec2vec has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability.
 GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017).
@@
 Our point of departure is still universal geometry, but we pursue a different approach.
-We draw inspiration from the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations.
+We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations.
 Unfortunately, stitching requires access to pairs of matching representations.
@@
 Moschella et al.
-(2023), inspired by Olah (2015), utilized these observations by proposing relative representations as a mathematically principled framework for comparing embedding spaces.
+(2023) operationalized the observations of Olah (2015) and proposed relative representations as a mathematically principled framework for comparing latent spaces.
 Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable.
-This work provided both theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
+This work provided both a theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
 Given a set of anchor points $\{\mathbf{a}_{i}\}_{i=1}^{k}$ in an embedding space, the relative representation of any point $\mathbf{x}$ is defined as:
@@
 Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content.
-[^1] More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$.
+More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$.
 We will not take advantage of this flexibility in the paper.
+[^1] Of course, $f_{1}(x)=f_{2}(x)=\text{const}$ would be a trivial solution, without additional constraints.
+So, similarity needs to be evaluated in comparison to other vectors in the dataset.
 Algorithm 3 Refine-1: Matching-Based Refinement
@@
 e5 0.98 1.11 (0.0) 0.28 (0.00) 0.52 (0.00) 0.53 (0.00) 0.99 (0.00) 1.05 (0.00)
+gte 0.95 1.18 (0.0) 0.35 (0.04) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.02 (0.00)
 gran. 0.99 1.02 (0.0) 0.31 (0.02) 0.58 (0.00) 0.57 (0.00) 1.00 (0.00) 1.01 (0.00)
@@
 e5 0.84 2.88 (0.2) 0.22 (0.04) 0.48 (0.00) 0.49 (0.00) 0.98 (0.00) 1.06 (0.00)
+$\text{gte}^{*}$ 0.93 2.31 (0.1) 0.29 (0.02) 0.50 (0.00) 0.51 (0.00) 0.98 (0.00) 1.05 (0.00)
+gran. 0.95 1.22 (0.0) 0.31 (0.01) 0.57 (0.00) 0.58 (0.00) 0.99 (0.00) 1.01 (0.00)
+gte $\text{gtr}^{*}$ 0.91 2.64 (0.1) 0.26 (0.02) 0.51 (0.00) 0.52 (0.00) 0.98 (0.00) 1.04 (0.00)
+stella 1.00 1.00 (0.0) 0.42 (0.02) 0.70 (0.00) 0.71 (0.00) 1.00 (0.00) 1.00 (0.00)
+e5 0.99 5.91 (0.5) 0.33 (0.05) 0.60 (0.00) 0.61 (0.00) 1.00 (0.00) 1.00 (0.00)
 gran. 0.98 1.08 (0.0) 0.37 (0.02) 0.58 (0.00) 0.59 (0.00) 0.99 (0.00) 1.02 (0.00)
@@
 e5 1.00 1.00 (0.0) 0.43 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
+gte 1.00 1.00 (0.0) 0.42 (0.00) 0.69 (0.00) 0.69 (0.00) 1.00 (0.00) 1.00 (0.00)
 gran. 0.99 2.20 (0.2) 0.30 (0.00) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.04 (0.00)
@@
 stel. 1.00 1.00 (0.0) 0.44 (0.01) 0.64 (0.00) 0.65 (0.00) 1.00 (0.00) 1.00 (0.00)
+gte 1.00 1.00 (0.0) 0.33 (0.00) 0.61 (0.00) 0.62 (0.00) 1.00 (0.00) 1.00 (0.00)
 Table 1: Results show mean $\pm$ standard deviation over 3 runs.
@@
 Bold and underline combined indicate better performance, but only within a $0.01$ margin.
+For the gtr-gte pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
 4 Methodology
@@
 In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$.
-Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
+Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
 This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces.
@@
 where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices.
-For solving QAP, we use $2$-OPT (running 30 times and choosing the one with the highest alignment score for further robustness).
+For solving QAP, we use $2$-OPT (Croes, 1958).
+We run it 30 times and choose the permutation with the highest alignment score.
+This helps improve the robustness of the alignment over a single run of 2-OPT.
+Due to the small number of cluster centroids, this takes only a few seconds.
 Relative Representations.
@@
 This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima.
-Pseudo-code is provided in Algorithm 3 & 4.
+Pseudo-code is provided in Algorithms 3 & 4.
 Refine-1: Matching-based Refinement.
@@
 $\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
-with $\alpha=0.5$, though the results appear robust to other choices as well.
+In the experiments below, we use $\alpha=0.5$.
 Refine-2: Clustering-based Refinement.
@@
 (2025).
-We run the algorithm for all pairs among four sets of text encoders used in Jha et al.
-(2025): stella, gtr, granite, and e5 (Zhang et al., 2025; Ni et al., 2021; Awasthy et al., 2025; Wang et al., 2024).
-Experiments use K-Means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT (Croes, 1958).
+We run the algorithm for all pairs among the 5 sets of text encoders used in Jha et al.
+(2025).
+The encoders are summarized in Table 2.
+Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT.
 We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$ and compute sentence embeddings under all possible encoders.
@@
 We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages.
-[^3] Cosine similarity in itself is indicative, but it can also be misleading.
-Cosine similarity in mini-vec2vec underestimates the similarity, in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
+Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading.
+Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
+Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry.
+Our setup does not change the geometry.
+[^3] Apart from translation and scaling, which, formally speaking, are a form of transition to a third space, but in a very restricted sense.
 Hyperparameters.
@@
 For exponential smoothing, we use $\alpha=0.5$.
-Note this is a rather economical choice of hyperparameters, which we would not have suggested to use a priori.
+Note this is a rather economical choice of hyperparameters, which we would not suggest using a priori.
 It turned out to work well, regardless.
 The full hyperparameter configuration used in the experiments is also documented in the code linked.
+For the gtr-gte pair, we increase the number of clusters to $c=30$.
+We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
 Results.
 In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table.
-Our advantage is more noticeable in the rank metric.
-Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already at the end of the first refinement step.
-Unlike vec2vec, our method does not collapse on the $e5$ and gtrgtr pair.
+The improvement is more noticeable in the rank metric.
+Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already by the end of the first refinement step.
+Unlike vec2vec, our method does not collapse on the $e5$-gtrgtr pair or the gtrgtr-gtegte pair.
 We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup.
@@
 Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise.
-Our results are not meant to show superiority, but rather matching performance for a fraction of the cost – except in the cases where vec2vec converges poorly, in which we observe mini-vec2vec’s superior behavior.
-Computationally, one run is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
+The superior results are an additional benefit, but even where the results are not superior, but rather matching performance, it is achieved for a fraction of the cost.
+Moreover, even where vec2vec collapses, we observe mini-vec2vec still reaches good performance.
+Computationally, one run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
 Analysis.
 We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations.
-This reproducible convergence behavior suggests the optimization landscape may contain well-connected regions or multiple basins with similar minima.
+This reproducible convergence behavior suggests the optimization landscape is well-structured.
 The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective.
@@
 The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
+Model Params Backbone Paper Year
+gtr 110M T5 Ni et al. 2021
+e5 109M BERT Wang et al. 2022
+gte 109M BERT Li et al. 2023
+stella 109M BERT Zhang et al. 2023
+granite 278M RoBERTa Awasthy et al. 2024
+Table 2: Embedding models used in our experiments
+6 Discussion
+Choice of Hyperparameters
+The mini-vec2vec algorithm is reasonably stable under different choices of parameters, often converging to similar stationary points.
+Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and frequently have a monotone behavior.
+For example, generally speaking, increasing the following hyperparameters should improve the outcome (or at least not harm).
+Importantly, while some of these trends have been observed empirically in some setups, a full experimental analysis has not been made, and so the explanations below on how to choose hyperparameters are suggestions at best, and should not be trusted blindly.
+  • Number of concatenated representations: The more concatenated alignments we use, the more likely it is for noisy assignments to be diluted, emphasizing the signal from true assignments.
+Note that this helps stabilize the algorithm, meaning it is less likely to meet failure modes in the average and worst cases.
+However, even with fewer alignments, good alignments can be reached, but the probability of failure modes might increase.
+  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows.
+This makes it a safe bet to increase the number of iterations if one is unsure whether convergence has been reached.
+Across all cases examined, we found convergence to be quite fast.
+Almost without exception, 100 iterations were more than enough, by a large margin.
+  • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise.
+The larger the sample size is, the closer this approximation is to the data.
+We generally consider higher sampling rates to be better.
+Other hyperparameters are less obvious.
+Number of clusters in the approximate matching step is likely good in moderation.
+While it gives us more landmark points for alignment, it can theoretically lead to instabilities, because the assignment of very granular centroids can be noisy.
+Increasing number of neighbors is often beneficial, to a point.
+It helps reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position.
+Note that this is in contrast to multilingual word embeddings, where essentially every point has a ground-truth translation, and it makes sense to rely on a single neighbor.
+Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance eventually.
+Finally, increasing the exponential smoothing parameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores.
+This is likely due to overfitting.
+The smoothing parameter is implicitly responsible for the maximum distance of the matrix from being orthogonal.
+While this (alleged) overfitting phenomenon has been observed by us empirically, we did not explore it rigorously, and it might not be a common case.
+Additionally, increasing $\alpha$ makes convergence slightly slower.
+Unlike many deep learning setups – GANs in particular – the range of hyperparameters that lead to strong performance is generously wide.
+Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter in order to improve effectiveness, if needed.
+Failure Modes
+We find that despite the many experiments we ran, including ones with different configurations, failure modes were extremely rare.
+In fact, we didn’t experience any failure modes except with the gte model.
+Moreover, all failure modes were around the gte-gtr pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in gtr $\rightarrow$ gte – they are just rarer.
+Crucially, failure modes are not catastrophic.
+The failure modes are characterized by worse top-1 accuracy and average rank (and slightly lower cosine similarity throughout the refinement steps.
+Often also noticeable in the initial step, but not always).
+But the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
 7 Related Work
@@
 (2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching.
+Similarly, Dar et al.
+(2023) have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors.
 Huh et al.
 (2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
-The task of usupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018).
+The task of unsupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018).
 Most of the works rely on training a GAN.
@@
 By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments.
-From a security perspective, efficient alignment methods raise important considerations about intellectual property protection in embedding spaces.
+From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces.
 If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al.
 (2025).
+Acknowledgements
+We thank John X. Morris for useful feedback. We thank Google for making CPU and GPU notebooks freely available.

### v3

@@
 Abstract
-We build upon vec2vec, a procedure designed to align text embedding spaces without parallel data.
-vec2vec finds a near-perfect alignment, but it is expensive and unstable.
+We build upon vec2vec, a procedure designed to align embedding spaces without parallel data.
+While it usually finds a near-perfect alignment, it is expensive and unstable.
 We present mini-vec2vec, a simple and efficient alternative that is highly robust and incurs a much lower computational cost.
-Moreover, the learned mapping is a linear transformation.
-Our method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement.
-Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding their results.
-The method’s stability and interpretable algorithmic steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
-Algorithm 1 Unsupervised Embedding Alignment
-1:Embedding matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$, $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$
-2:Number of clusters $c$, neighbors $k$, iterations $T$, smoothing parameter $\alpha$, runs $s$, clusters $c^{\prime}$, neighbors $k^{\prime}$, sample size $n_{s}$
-3:Linear transformation matrix $\mathbf{W}$
-4:Normalize embeddings: $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}\leftarrow\text{center and normalize to unit sphere}$
-5:Create pseudo-pairs: $\mathcal{C}\leftarrow\operatorname{AnchorAlignment}(\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B},c,k,s)$ $\triangleright$ See Alg. 2
-6:Estimate initial mapping: $\mathbf{W}\leftarrow\text{Procrustes analysis on }\mathcal{C}$
-7:Refine-1: Apply matching-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 3
-8:Refine-2: Apply clustering-based refinement on $\mathbf{W}$ $\triangleright$ See Alg. 4
-9:return $\mathbf{W}$
+Moreover, the transformation that mini-vec2vec learns is linear.
+The method consists of three main stages: a tentative matching of pseudo-parallel embedding vectors, transformation fitting, and iterative refinement.
+Our linear alternative exceeds the original instantiation of vec2vec by orders of magnitude in efficiency, while matching or exceeding its alignment performance.
+The method’s stability and interpretable steps facilitate scaling and unlock new opportunities for adoption in new domains and fields.
 1 Introduction
-Representation learning has revolutionized natural language processing by producing embedding models that capture rich semantic structure in high-dimensional vector spaces.
-These learned representations encode complex linguistic relationships, enabling downstream tasks such as semantic similarity and information retrieval.
-However, different models, training procedures, or even different random initializations of the same architecture can produce embedding spaces that, while semantically meaningful, are not directly comparable due to arbitrary rotations, reflections, and translations in their coordinate systems.
-The problem of aligning embedding spaces without parallel supervision – that is, without access to pairs of embeddings that represent identical semantic content – has emerged as a fundamental challenge with far-reaching implications.
-Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (Conneau et al., 2018; Zhang et al., 2017), cross-modal alignment between text and image representations (Maniparambil et al., 2024), and privacy concerns (Song and Raghunathan, 2020; Morris et al., 2023).
-The recent vec2vec method (Jha et al., 2025) demonstrated that unsupervised alignment is indeed possible using CycleGAN (Zhu et al., 2020), an adversarial training framework.
-Their approach is theoretically grounded in the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces.
-Under this hypothesis, semantic relationships are preserved across models in terms of relative distances and angles, suggesting that alignment should be achievable through geometric transformations.
-While vec2vec has proven effective, adversarial training frameworks bring inherent challenges that limit their practical applicability.
-GANs are notoriously computationally intensive, requiring careful hyperparameter tuning and often exhibiting training instability characterized by mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017).
-These methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
-Our point of departure is still universal geometry, but we pursue a different approach.
-We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (Lenc and Vedaldi, 2015; Bansal et al., 2021), which have demonstrated that simple affine transformations often suffice to align representations.
-Unfortunately, stitching requires access to pairs of matching representations.
-Our key methodological insight is that with universal geometry, structural similarities should be detectable and exploitable, even without adversarial training.
-Specifically, the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts should provide sufficient signal for alignment, even in the complete absence of parallel supervision.
-To this end, we utilize the Relative Representation framework of Moschella et al.
-(2023), which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
-Our contribution is mini-vec2vec, a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (running on commodity CPU hardware), training stability (deterministic components with controllable stochastic elements), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples).
+Sentence embeddings capture rich semantic structure in high-dimensional vector spaces.
+They encode complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval.
+However, different architectures, training procedures, and even different models with the same architecture produce embedding spaces that cannot be compared directly, because they use a different coordinate system, a different “internal language”.
+The problem of aligning embedding spaces without parallel supervision – that is, without pairs of matching encodings of the same data – has emerged as a fundamental challenge with far-reaching implications.
+Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (conneau2017word; zhang-etal-2017-adversarial), cross-modal alignment between text and image representations (maniparambil2024; schnaus2025), and privacy concerns (song2020informationleakage; text2vec).
+The recent vec2vec method (jha2025vec2vec) has demonstrated that unsupervised alignment is indeed possible using CycleGAN (zhu2017cyclegan).
+Their approach is theoretically grounded in a principle called the Universal Geometry or the Platonic Representation Hypothesis (huh2024platonic), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces, even if their coordinate systems differ.
+This means that there is a way to rotate the coordinate systems so they “understand” each other – and vec2vec does precisely that; it identifies a rotation that makes the coordinate systems aligned.
+While vec2vec has proven effective, adversarial training frameworks like CycleGAN bring inherent challenges that limit their practical applicability.
+Training them is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (saxena2023gan_survey; goodfellow2017nips2016).
+Moreover, they typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
+Our point of departure is still Universal Geometry, but we take a different approach.
+We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (lenc2014stitching; bansal2021revisiting), which have demonstrated that simple affine transformations often suffice to align representations.
+But unfortunately, stitching assumes access to pairs of matching representations.
+Motivated by the Platonic Representation Hypothesis, we hypothesize that we can nonetheless uncover this linear alignment without even needing adversarial training.
+Specifically, we hypothesize that the similarity in the geometric structure – the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts – should provide sufficient signal for alignment without parallel supervision and without need for adversarial methods.
+To this end, we utilize the Relative Representations framework of moschella2022relative, which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
+We then use it as a stepping stone to produce a noisy alignment that we refine with further steps.
+[^1] In Section 6, we discuss a variant that sidesteps the use of relative representations, but we tested it far less.
+mini-vec2vec is a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (very fast and running on commodity CPU hardware), training stability (good performance and robustness to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples).
 Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
@@
 The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales.
-Early work by Olah (2015) revealed striking structural regularities in representations of different convolutional networks, showing distances between object representations are roughly the same regardless of architecture.
-These findings suggested that the geometric organization of learned features might be more universal than previously assumed.
-The Platonic Representation Hypothesis, formally articulated by Huh et al.
-(2024), represents the most comprehensive theoretical framework for understanding universal geometry.
+Early work by colah2015visualizing has revealed striking structural similarities in representations of different convolutional networks, [^2] Olah briefly refers to even earlier ideas (e.g., a concept called “meta-tSNE” by Hinton), but we failed to locate them.
+suggesting that the geometric organization of learned features might be more universal than previously assumed.
+The Platonic Representation Hypothesis, formally articulated by huh2024platonic, represents the most comprehensive theoretical framework for understanding Universal Geometry.
 The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain.
-Moschella et al.
-(2023) operationalized the observations of Olah (2015) and proposed relative representations as a mathematically principled framework for comparing latent spaces.
-Their key insight was that while absolute coordinates may vary arbitrarily across models, relative similarity structures – captured through pairwise distances to a shared set of anchor data points – remain remarkably stable.
-This work provided both a theoretical foundation and empirical evidence for the existence of model-agnostic geometric properties.
-Given a set of anchor points $\{\mathbf{a}_{i}\}_{i=1}^{k}$ in an embedding space, the relative representation of any point $\mathbf{x}$ is defined as:
-$\mathbf{r}(\mathbf{x})=[f(\mathbf{x},\mathbf{a}_{1}),f(\mathbf{x},\mathbf{a}_{2}),\ldots,f(\mathbf{x},\mathbf{a}_{k})]^{T},$
-where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity). Two representations from different models will have close relative representations if they represent similar objects.
-Algorithm 2 Algorithm for creating pseudo-pairs using noisy anchors
-1:Normalized embeddings $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, clusters $c$, neighbors $k$, runs $s$
-2:Set of pseudo-parallel pairs $\mathcal{C}$
-3:Initialize collections: $\mathcal{R}_{A}\leftarrow[],\mathcal{R}_{B}\leftarrow[]$
-4:for $i=1$ to $s$ do $\triangleright$ Multiple runs for robustness
-5:  Cluster both spaces: $\{\mathbf{c}^{A}_{i,j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{A})$, $\{\mathbf{c}^{B}_{i,j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{B})$
-6:  Compute similarities: $\mathbf{S}^{A},\mathbf{S}^{B}\leftarrow\text{cosine similarities between centroids}$
-7:  Find correspondence: $\mathbf{P}^{*}_{i}\leftarrow\text{solve QAP}(\mathbf{S}^{A},\mathbf{S}^{B})$ using multiple 2-OPT runs
-8:  Build relative representations: $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}\leftarrow\text{cosine similarities to anchors}$
-9:  Store: append $\mathbf{r}^{A}_{i},\mathbf{r}^{B}_{i}$ to $\mathcal{R}_{A},\mathcal{R}_{B}$
-10:end for
-11:Concatenate: $\mathbf{r}^{A}\leftarrow\text{concatenate all runs in }\mathcal{R}_{A}$
-12:Concatenate: $\mathbf{r}^{B}\leftarrow\text{concatenate all runs in }\mathcal{R}_{B}$
-[…]
-14:Create pseudo-pairs: $\mathcal{C}\leftarrow\left\{(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{j}\mathbf{x}^{B}_{j}):\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})\right\}$
-15:return $\mathcal{C}$
+Having the same geometry implies that the distance between objects should be independent of the models themselves.
+For example, if one wants to compute the semantic similarity of “I love your dog” and “I love my cat” using cosine similarity between their embeddings, the answer should be roughly the same regardless of the underlying embedding model.
+Intuitively, this empirical observation stands to reason, as semantic similarity originates from the semantic meaning that comes from the data itself.
+Below, we discuss the framework of moschella2022relative that extends this observation into a powerful tool.
+Relative Representations
+moschella2022relative have operationalized the observations of colah2015visualizing and proposed to use relative representations as a mathematically principled framework for comparing latent spaces.
+While absolute coordinates vary arbitrarily across models, we know that relative information, such as similarity between two objects, is (approximately) agnostic to the choice of embedding model.
+They propose to use a shared set of anchor points that other data points will be compared to as a way to ground the position of data points in a model-agnostic way.
+Denote the space of data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality) and the embedding function by $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$.
+Given a set of anchor data points $\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under $E$ is:
+$\mathbf{r}_{E}(\mathbf{z})=[f(E(\mathbf{z)},E(\mathbf{a}_{1}));f(E(\mathbf{z}),E(\mathbf{a}_{2}));\ldots;f(E(\mathbf{z}),E(\mathbf{a}_{k}))],$
+where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity) and ; represents concatenation.
+The power of relative representations is that, as shown by colah2015visualizing and moschella2022relative, for two encoders $E_{1},E_{2}$, the relative representations of the same data point is roughly the same, i.e., $\mathbf{r}_{E_{1}}(\mathbf{z})\approx\mathbf{r}_{E_{2}}(\mathbf{z})$.
+In other words, the function $\mathbf{r}_{E}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ is itself an embedding that maps a data point into $k$-dimensional space in a way that is approximately independent of $E$ (for a sufficiently well-trained $E$) and only depends on the choice of anchor points.
 3 Problem Formulation
@@
 Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points.
-This approach offers key advantages:
+This approach offers several advantages:
   • Generalization: New embeddings can be translated without re-applying the algorithm
   • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
-  • Out-of-distribution capability: The parameterized nature enables generalization beyond the training data
-The universal geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries where cosine similarities remain approximately invariant across embedding spaces.
-These geometric similarities enable the identification of this alignment.
+  • Out-of-distribution capability: The parameterized nature can (potentially) enable generalization beyond the training distribution
+The Universal Geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries – relative similarities between data points remain approximately the same regardless of the embedding model.
 Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$.
-These embeddings represent semantic content from potentially different models, architectures, or training procedures, but we assume they encode similar semantic information from the same underlying distribution $\mathcal{D}$.
-Unlike more traditional setups, there does not exist a pairing between data points, i.e., there is no overlap between the sentences themselves.
+The sentences come from the same underlying distribution $\mathcal{D}$, but – unlike most other works – there is no overlap between the sentences themselves – only the distribution is shared.
 We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector.
-The fundamental challenge is that while both spaces may contain similar semantic structures, they are related by an unknown transformation.
-Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(\mathbf{x}^{A}_{i})\approx\mathbf{x}^{B}_{j}$ when $\mathbf{x}^{A}_{i}$ and $\mathbf{x}^{B}_{j}$ represent similar semantic content.
-More generally, the setup allows for learning a pair of functions $f_{1},f_{2}$ sending embeddings to a third space, such that $f_{1}(\mathbf{x}^{A}_{i})\approx f_{2}(\mathbf{x}^{B}_{j})$.
-We will not take advantage of this flexibility in the paper.
-[^1] Of course, $f_{1}(x)=f_{2}(x)=\text{const}$ would be a trivial solution, without additional constraints.
-So, similarity needs to be evaluated in comparison to other vectors in the dataset.
-Algorithm 3 Refine-1: Matching-Based Refinement
-1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
-2:Iterations $T$, smoothing parameter $\alpha$, neighbors $k^{\prime}$, sample size $n_{s}$
-3:Refined transformation matrix $\mathbf{W}$ after matching-based refinement
-4:for $t=1$ to $T$ do
-5:  Sample embeddings: $\mathbf{X}_{\text{sample}}\leftarrow\text{random sample of }n_{s}\text{ rows from }\hat{\mathbf{X}}_{A}$
-6:  Transform samples: $\mathbf{X}_{\text{transformed}}\leftarrow\mathbf{X}_{\text{sample}}\mathbf{W}$
+The fundamental challenge is that while both spaces may contain similar geometric structures, they are related by an unknown transformation.
+Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a datapoint $\mathbf{z}\in\mathcal{Z}$.
+More generally, the setup described in jha2025vec2vec allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$.
+We will not take advantage of this flexibility in this paper.
+[^3] It also complicates the analysis of the results.
+For example, $f_{A}=f_{B}=\text{const}$ would be a trivial solution if we don’t impose additional constraints.
+Similarity then needs to be assessed more carefully.
+4 Methodology
+4.1 Motivation
+In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$.
+Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
+This is motivated by the Universal Geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces.
+Later on, we will relax this assumption and ultimately the matrix will be an exponentially-smoothed average of orthogonal matrices, enforcing orthogonality softly.
+Our linear approach, mini-vec2vec, addresses key limitations of the current approach: GANs are notoriously unstable and hard to train, demand extensive computational resources and training time.
+They often require tedious hyperparameter tuning and sometimes multiple re-runs to converge to good solutions.
+On the other hand, we know that in supervised cases (model stitching), there exists an affine mapping between the latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently.
+mini-vec2vec achieves that, and it has several appealing properties: it requires significantly less data than vec2vec (60k vs.
+1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
+Our method resembles the approach described in hoshen-wolf-2018-non for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement.
+This paradigm is very powerful, and we adopt a similar approach.
+Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because there is no small exhaustive set of objects (words) that need to be matched.
+Moreover, the very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small.
+Therefore, word embedding alignment is rarely considered in the non-overlapping datasets setup (even in multilingual matching, words mostly have a ground-truth translation).
+We find that applying the algorithm from hoshen-wolf-2018-non naively to our use case doesn’t work.
+The approximate matching stage, in particular, requires more delicate treatment in our case.
+The iterative refinement step is also improved by a slightly more subtle approach.
+def mini_vec2vec(X_A, X_B):
+X_A, X_B = normalize(X_A), normalize(X_B)
+C = anchor_pairs(X_A, X_B)
+W = procrustes(C)
+W = refine_1(X_A, X_B, W)
+W = refine_2(X_A, X_B, W)
+return W
+def anchor_pairs(X_A, X_B):
+R_A_list, R_B_list = [], []
+for i in range(num_runs):
+C_A = k_means(X_A, num_clusters)
+C_B = k_means(X_B, num_clusters)
+S_A = similarity_function(C_A, C_A)
+S_B = similarity_function(C_B, C_B)
+pi = solve_QAP(S_A, S_B)
+C_B_permuted = permute(C_B, pi)
+R_A_list.append(similarity_function(X_A, C_A))
+R_B_list.append(similarity_function(X_B, C_B_permuted))
+R_A, R_B = concatenate(R_A_list), concatenate(R_B_list)
+neighbors = knn(R_A, R_B, num_neighbors=k)
+X_matched = average_nearest_neighbors(X_B, neighbors)
+return X_A, X_matched
+Listing 1: Pseudo-code for mini-vec2vec (refinement algorithms deferred to Listing LABEL:lst:refine)
+We will now describe our algorithm in detail. Pseudo-code is provided in Listing LABEL:lst:full_algo.
+4.2 Preprocessing
+We first normalize the embeddings in both spaces. We center each set around its respective mean and normalize the embeddings to the unit hypersphere:
+$\boldsymbol{\mu}_{A}$ $=\frac{1}{n_{A}}\mathbf{X}_{A}^{T}\mathbf{1}_{n_{A}},\quad\tilde{\mathbf{X}}_{A}=\mathbf{X}_{A}-\mathbf{1}_{n_{A}}\boldsymbol{\mu}_{A}^{T}$
+$\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
+where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
+4.3 Approximate Matching
+To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching.
+We will first describe a single run, and then explain how we robustify it by ensembling multiple runs:
+Centroid Matching
+The reason we choose to align cluster centroids is that they roughly represent the central landmarks in the dataset.
+Since the data comes from a shared distribution $\mathcal{D}$, we expect clusters to be similar overall, and all that remains to do is to find the alignment between them.
+Of course, clusters are not very consistent even with exactly the same data, so the approach needs some tolerance to errors (hence the robustification step below).
+We perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$.
+We then compute similarity matrices between the centroids independently in each embedding space:
+$S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
+where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
+We then find the optimal matching between cluster centroids from different embedding spaces by solving the Quadratic Assignment Problem (QAP):
+$\pi^{*}=\arg\max_{\pi\in\Pi_{k}}\;\sum_{i=1}^{k}\sum_{j=1}^{k}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
+where $\pi:\{1,2,...,k\}\rightarrow\{1,2,...,k\}$ is a permutation from the set $\Pi_{k}$ of all permutations on $k$ elements.
+Due to the small number of cluster centroids, this takes only a few seconds.
+Relative Representations
+Once we have identified an alignment between $A\text{-}$ and $B\text{-centroids}$, we use the aligned centroids as anchors.
+We represent each embedding through its relationships to anchor points.
+For embedding $\mathbf{x}^{A}_{i}$, we construct:
+$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]$
+Similarly, for $\mathbf{x}^{B}_{i}$, we compute (note that we need to re-arrange the indices with $\pi$):
 […]
-8:  Average neighbors: $\mathbf{X}_{\text{matched}}\leftarrow\text{mean of neighbors for each sample}$
-9:  Construct pseudo-pairs: $\mathcal{C}^{\prime}\leftarrow\{(\mathbf{X}_{\text{sample}},\mathbf{X}_{\text{matched}})\}$
-10:  Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
-11:  Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
-12:end for
-Algorithm 4 Refine-2: Clustering-Based Refinement
-1:Embedding matrices $\hat{\mathbf{X}}_{A},\hat{\mathbf{X}}_{B}$, initial transformation $\mathbf{W}$
-2:Smoothing parameter $\alpha$, number of clusters $c^{\prime}$
-3:Refined transformation matrix $\mathbf{W}$ after clustering-based refinement
-4:Cluster space A: $\{\mathbf{c}^{A}_{j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{A},c^{\prime})$
-5:Transform centroids: $\{\mathbf{c}^{A\to B}_{j}\}\leftarrow\{\mathbf{c}^{A}_{j}\mathbf{W}\}$
-6:Cluster space B with seeds: $\{\mathbf{c}^{B}_{j}\}\leftarrow\text{k-means}(\hat{\mathbf{X}}_{B},c^{\prime},\text{init}=\{\mathbf{c}^{A\to B}_{j}\})$
-7:Construct pairs $\mathcal{C}^{\prime}$ by matching cluster centroids
-8:Estimate new mapping: $\mathbf{W}_{\text{new}}\leftarrow\text{Procrustes analysis on }\mathcal{C}^{\prime}$
-9:Update with smoothing: $\mathbf{W}\leftarrow(1-\alpha)\mathbf{W}+\alpha\mathbf{W}_{\text{new}}$
-10:return $\mathbf{W}$
+Robustification via Ensembling
+We improve robustness by repeating the process described above and concatenating relative representations from multiple runs:
+$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk},\qquad\text{similarly for space }B$
+This concatenation creates richer structural signatures that are more robust to noise that comes from bad alignments.
+4.4 Mapping Estimation
+We construct pseudo-parallel pairs by sending each element in space $A$ to the average of its $k$ nearest neighbors from space $B$ (based on similarity in relative space):
+$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\hskip-3.00003pt\sum_{\tiny\hskip 2.94443pt{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\hskip-13.25018pt}\hskip-3.00003pt\mathbf{x}^{B}_{j}\,\,\,\,\right):\,i=1,\ldots,n_{A}\right\}$
+where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space.
+Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$.
+Optimal orthogonal transformation is obtained by Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
+4.5 Iterative Refinement
+The initial transformation provides a coarse alignment that we refine with two complementary strategies. Pseudo-code is provided in Listing LABEL:lst:refine.
+Refine-1: Matching-based Refinement
+This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.4), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space.
+The method is cheap and it runs for many iterations (50–100), using subsampling to boost computational efficiency further.
+We use exponential smoothing for transformation updates:
+$\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
+Refine-2: Clustering-based Refinement
+This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids.
+We begin by clustering the embeddings in space $A$.
+We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization.
+The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors.
+The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids.
+This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$.
+Similar to the first strategy, we use exponential smoothing.
+Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1.
+Two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
+def refine_1(X_A, X_B, W):
+for t in range(num_iters):
+X_sample = sample(X_A, subsampling_rate)
+X_transformed = X_sample @ W
+neighbors = knn(X_transformed, X_B, num_neighbors=k2)
+X_sample_matched = average_nearest_neighbors(X_B, neighbors)
+W_new = procrustes(X_sample, X_sample_matched)
+W = (1 - alpha) * W + alpha * W_new
+return W
+def refine_2(X_A, X_B, W):
+C_A = k_means(X_A, num_clusters2)
+C_B = k_means(X_B, num_clusters2, init=C_A @ W)
+W_new = procrustes(C_A, C_B)
+return (1 - alpha) * W + alpha * W_new
+Listing 2: Pseudo-code for the iterative refinement step
+5 Experiments
+We replicate the main experiment from jha2025vec2vec.
+We run our algorithm for all pairs of encoders used in jha2025vec2vec.
+Details about the encoders are provided in Table 2.
+Experiments use k-means clustering from scikit-learn (scikit-learn) and QAP is solved with the scipy (2020SciPy-NMeth) implementation of 2-OPT.
+We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score.
+This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
+We take a random subset of the Natural Questions dataset (natural_questions) of size $60,000$ and compute sentence embeddings under all possible encoders.
+We leave $8192$ sentences for evaluation.
+For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences.
+All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate consistency.
+We compare our results to the numbers reported in jha2025vec2vec.
+Metrics
+We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by:
+$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\text{NN}(\mathbf{W}\mathbf{a}_{i})=\mathbf{b}_{i}),$
+where $\delta$ is the indicator function, $\text{NN}(\cdot)$ is the nearest neighbor of source embedding $\mathbf{a}_{i}$ in the target space, and $\mathbf{b}_{i}$ is its true corresponding target embedding.
+The second is the Average Rank, the mean position of the true match in the sorted list of distances from the query to all targets,
+$\text{AvgRank}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}(\mathbf{b}_{i}|\mathbf{a}_{i}).$
+We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages.
+Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading.
+Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
+Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry.
+Our setup does not change the geometry (apart from aparametric translation and scaling in the preprocessing step).
+Hyperparameters
+We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range.
+We make an economical choice of hyperparameter values due to the large number of runs, but we find the choice to be sufficiently robust still.
+We use $s=30$ runs of relative representation alignment with $c=20$ clusters.
+In Mapping Estimation, we use $k=50$ neighbors.
+We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters.
+For exponential smoothing, we use $\alpha=0.5$.
+The full hyperparameter configuration used in the experiments is also documented in the code linked.
+For the $\textit{gtr}\leftrightarrow\textit{gte}$ pair, we increase the number of clusters to $c=30$.
+We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
 vec2vec mini-vec2vec
@@
 e5 0.98 1.11 (0.0) 0.28 (0.00) 0.52 (0.00) 0.53 (0.00) 0.99 (0.00) 1.05 (0.00)
-gte 0.95 1.18 (0.0) 0.35 (0.04) 0.54 (0.00) 0.56 (0.00) 0.99 (0.00) 1.02 (0.00)
+gte 0.95 1.18 (0.0) 0.35 (0.04) 0.54 (0.00) 0.56 (0.00) 0.98 (0.00) 1.05 (0.00)
 gran. 0.99 1.02 (0.0) 0.31 (0.02) 0.58 (0.00) 0.57 (0.00) 1.00 (0.00) 1.01 (0.00)
@@
 $\text{gte}^{*}$ 0.93 2.31 (0.1) 0.29 (0.02) 0.50 (0.00) 0.51 (0.00) 0.98 (0.00) 1.05 (0.00)
-gran. 0.95 1.22 (0.0) 0.31 (0.01) 0.57 (0.00) 0.58 (0.00) 0.99 (0.00) 1.01 (0.00)
+gran. 0.95 1.22 (0.0) 0.31 (0.01) 0.57 (0.00) 0.58 (0.00) 0.99 (0.00) 1.03 (0.00)
 gte $\text{gtr}^{*}$ 0.91 2.64 (0.1) 0.26 (0.02) 0.51 (0.00) 0.52 (0.00) 0.98 (0.00) 1.04 (0.00)
@@
 Bold and underline combined indicate better performance, but only within a $0.01$ margin.
-For the gtr-gte pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
-4 Methodology
-4.1 Motivation
-In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$.
-Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
-This is motivated by the universal geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces.
-Later on, we will relax this assumption.
-Ultimately, the matrix will be the exponentially-smoothed average of a sequence of orthogonal matrices, enforcing orthogonality softly.
-Our linear approach addresses key limitations of existing methods: GANs are notoriously unstable and hard to train with many hyperparameters, require extensive computational resources and training time, and often need multiple seeds to achieve good solutions.
-In contrast, we know that in supervised cases (model stitching), there usually exists an affine mapping between latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently.
-Our method requires significantly less data (60k vs.
-1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
-Our method resembles Hoshen and Wolf (2018)’s approach for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement.
-This scheme is very powerful, and we apply a similar method.
-We find that the approximate matching stage especially requires more delicate treatment in our case, so their approach is not applicable as-is.
-The stages incrementally improve alignment quality – each stage builds upon structural correspondences to establish increasingly accurate alignments:
-  • Approximate Matching: Find approximate pairs of embeddings from both spaces using landmark points as anchors for relative representations
-  • Mapping Estimation: Learn a linear orthogonal mapping from approximate matchings, where linearity and orthogonality provide an inductive bias to smooth out incorrect pairings
-  • Iterative Refinement: Use the initial solution as a stepping stone for better matching/mapping, iterating until convergence or budget exhaustion
-The overall process is described in Algorithm 1.
-4.2 Approximate Matching
-Preprocessing.
-We first address arbitrary translations and scalings across embedding spaces. Both spaces are centered around their respective means and normalized to the unit hypersphere:
-$\boldsymbol{\mu}_{A}$ $=\frac{1}{n_{A}}\mathbf{X}_{A}^{T}\mathbf{1}_{n_{A}},\quad\tilde{\mathbf{X}}_{A}=\mathbf{X}_{A}-\mathbf{1}_{n_{A}}\boldsymbol{\mu}_{A}^{T}$
-$\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
-where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
-Centroid Matching.
-To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching.
-Each run consists of k-means clustering followed by solving a quadratic assignment problem (QAP) for cluster correspondence.
-In each run, we first perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$.
-We then compute structural similarity matrices between cluster centroids:
-$S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
-where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
-We find the optimal matching between cluster centroids by solving the quadratic assignment problem:
-$\mathbf{P}^{*}=\operatorname{argmax}_{\mathbf{P}\in\Pi_{k}}\text{Tr}\left(\mathbf{S}_{A}^{T}\mathbf{P}\mathbf{S}_{B}\mathbf{P}^{T}\right)$
-where $\mathbf{P}$ is a permutation matrix from the set $\Pi_{k}$ of all $k\times k$ permutation matrices.
-For solving QAP, we use $2$-OPT (Croes, 1958).
-We run it 30 times and choose the permutation with the highest alignment score.
-This helps improve the robustness of the alignment over a single run of 2-OPT.
-Due to the small number of cluster centroids, this takes only a few seconds.
-Relative Representations.
-Once we have identified an alignment between centroids, we use the aligned centroids as anchors.
-For each run within the anchor discovery phase, we represent each embedding through its relationships to anchor points.
-For embedding $\mathbf{x}^{A}_{i}$ in a given run, we construct:
-$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]^{T}$
-where $\mathbf{r}^{A}_{i}$ is the relative representation of the $i$-th embedding in space $A$ with respect to the $k$ cluster centroids from that run.
-Robustification through Ensembling.
-We improve robustness by concatenating relative representations from multiple runs:
-$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk}$
-This concatenation creates richer structural signatures that are more robust to clustering initialization randomness.
-Anchor alignment algorithm is provided in Algorithm 2.
-4.3 Mapping Estimation
-We construct pseudo-parallel pairs by matching embeddings by similarity in relative space, averaging the $k$ nearest neighbors to reduce noise:
-$\mathcal{C}=\left\{\left(\mathbf{x}^{A}_{i},\frac{1}{k}\sum_{\mathbf{r}^{B}_{j}\in\mathcal{N}_{k}(\mathbf{r}^{A}_{i})}\mathbf{x}^{B}_{j}\right):i=1,\ldots,n_{A}\right\}$
-where $\mathcal{N}_{k}(\mathbf{r}^{A}_{i})$ denotes the $k$ nearest neighbors to $\mathbf{r}^{A}_{i}$ in the concatenated relative representation space.
-Note that the neighborhood information is obtained from the relative space (which is shared) and averaging takes place in the absolute space $B$.
-The orthogonal transformation is obtained via Procrustes analysis: $\mathbf{W}^{*}=\mathbf{V}\mathbf{U}^{T}$ where $\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}=\text{SVD}(\mathbf{A}^{T}\mathbf{B})$, and $\mathbf{A}$, $\mathbf{B}$ are matrices formed from the pseudo-parallel pairs in $\mathcal{C}$.
-4.4 Iterative Refinement
-The initial transformation provides a coarse alignment that we refine with two complementary strategies.
-This refinement is crucial to improve alignment quality, with Refine-2 helping Refine-1 escape local minima.
-Pseudo-code is provided in Algorithms 3 & 4.
-Refine-1: Matching-based Refinement.
-This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, and averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.3), with the crucial difference that now the similarity is performed in the ambient space $B$ rather than relative space.
-The method is cheap and it runs for many iterations (50–100) using subsampling to boost computational efficiency.
-We use exponential smoothing for transformation updates:
-$\mathbf{W}^{(t+1)}=(1-\alpha)\mathbf{W}^{(t)}+\alpha\mathbf{W}_{\text{new}}^{(t)}$
-In the experiments below, we use $\alpha=0.5$.
-Refine-2: Clustering-based Refinement.
-This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids.
-We begin by clustering the embeddings in space $A$.
-We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization.
-The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors.
-The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids.
-This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$.
-Similar to the first strategy, we use exponential smoothing.
-[^2] Note that the above justification for Refine-2 is our a priori rationale for using this method, and we do not make any claims as to whether this is the true reason for the ability of this method to improve the alignment.
-Our only claim is that we observe this step improves the alignment consistently, and despite the fact that the Refine-1 has already converged and no longer shows improvement.
-Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1.
-Curiously, two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
-5 Experiments
-We replicate the in-distribution experiment from Jha et al.
-(2025).
-We run the algorithm for all pairs among the 5 sets of text encoders used in Jha et al.
-(2025).
-The encoders are summarized in Table 2.
-Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011), QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT.
-We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$ and compute sentence embeddings under all possible encoders.
-We leave $8192$ sentences for evaluation.
-For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences.
-All experiments are run on a Colab notebook with a CPU runtime, and repeated three times for each pair to demonstrate consistency.
-We compare the results to the numbers reported in Jha et al.
-(2025).
-Metrics.
-We evaluate alignment quality using two primary metrics. The first is Top-1 Accuracy, the fraction of queries where the nearest neighbor in the target space is the correct match, defined by
-$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\delta(\text{NN}(\mathbf{W}\mathbf{a}_{i})=\mathbf{b}_{i}),$
-where $\delta$ is the indicator function, $\text{NN}(\cdot)$ is the nearest neighbor of source embedding $\mathbf{a}_{i}$ in the target space, and $\mathbf{b}_{i}$ is its true corresponding target embedding.
-The second is the Average Rank, the mean position of the true match in the sorted list of distances from the query to all targets,
-$\text{AvgRank}=\frac{1}{N}\sum_{i=1}^{N}\text{rank}(\mathbf{b}_{i}|\mathbf{a}_{i}).$
-We also report the mean cosine similarity after each step of the algorithm to show its progression over the stages.
-Cosine similarity in itself can be used as an evaluation metric, but it can also be misleading.
-Cosine similarity in mini-vec2vec underestimates the similarity in the original space, due to the removal of the mean vector, which is a dominant part of all vectors, accounting for $\sim 30\%-70\%$ of the vector’s norm.
-Also, setups that move the vectors to a third space, like vec2vec, have an advantage as they can change the geometry.
-Our setup does not change the geometry.
-[^3] Apart from translation and scaling, which, formally speaking, are a form of transition to a third space, but in a very restricted sense.
-Hyperparameters.
-We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range.
-We use $s=30$ runs of relative representation alignment with $c=20$ clusters.
-We use $k=50$ neighbors in Mapping Estimation.
-We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters.
-For exponential smoothing, we use $\alpha=0.5$.
-Note this is a rather economical choice of hyperparameters, which we would not suggest using a priori.
-It turned out to work well, regardless.
-The full hyperparameter configuration used in the experiments is also documented in the code linked.
-For the gtr-gte pair, we increase the number of clusters to $c=30$.
-We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
-Results.
+For the $gtr\leftrightarrow gte$ pair, we modify the number of clusters to $c=30$ in the approximate matching stage.
+Results
 In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table.
-The improvement is more noticeable in the rank metric.
-Importantly, the method is extremely robust, as is evidenced by the small standard deviation, already by the end of the first refinement step.
-Unlike vec2vec, our method does not collapse on the $e5$-gtrgtr pair or the gtrgtr-gtegte pair.
-We underline scores that are within a $0.01$ margin, as they might result from rounding noise and stochasticity, and a slightly different experimental setup.
+The improvement is most noticeable in the rank metric.
+Importantly, the method is extremely robust as is evidenced by the small standard deviation, already by the end of the first refinement step.
+Unlike vec2vec, our method does not collapse on the $\textit{e5}\leftrightarrow\textit{gtr}$ pair or the $\textit{gtr}\leftrightarrow\textit{gte}$ pair.
+We underline scores that are within a $0.01$ margin, as they might result from rounding noise, stochasticity, and a slightly different experimental setup.
 [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets.
-Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is a good compromise.
-The superior results are an additional benefit, but even where the results are not superior, but rather matching performance, it is achieved for a fraction of the cost.
-Moreover, even where vec2vec collapses, we observe mini-vec2vec still reaches good performance.
-Computationally, one run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
-Analysis.
-We observe that despite variation in the approximate matching phase (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations.
+Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is good enough.
+Analysis
+The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that it is achieved for a fraction of the cost.
+One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
+Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance.
+Analyzing the convergence of our algorithm, we observe that despite variation in the approximate matching step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations.
 This reproducible convergence behavior suggests the optimization landscape is well-structured.
@@
 However, applying Refine-2 successfully improves upon this converged solution.
-The two algorithms appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
+The two methods appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
 Model Params Backbone Paper Year
-gtr 110M T5 Ni et al. 2021
-e5 109M BERT Wang et al. 2022
-gte 109M BERT Li et al. 2023
-stella 109M BERT Zhang et al. 2023
-granite 278M RoBERTa Awasthy et al. 2024
+gtr 110M T5 gtr 2021
+e5 109M BERT e5 2022
+gte 109M BERT gte 2023
+stella 109M BERT stella 2023
+granite 278M RoBERTa granite 2024
 Table 2: Embedding models used in our experiments
 6 Discussion
-Choice of Hyperparameters
-The mini-vec2vec algorithm is reasonably stable under different choices of parameters, often converging to similar stationary points.
-Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and frequently have a monotone behavior.
-For example, generally speaking, increasing the following hyperparameters should improve the outcome (or at least not harm).
-Importantly, while some of these trends have been observed empirically in some setups, a full experimental analysis has not been made, and so the explanations below on how to choose hyperparameters are suggestions at best, and should not be trusted blindly.
-  • Number of concatenated representations: The more concatenated alignments we use, the more likely it is for noisy assignments to be diluted, emphasizing the signal from true assignments.
-Note that this helps stabilize the algorithm, meaning it is less likely to meet failure modes in the average and worst cases.
-However, even with fewer alignments, good alignments can be reached, but the probability of failure modes might increase.
-  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows.
-This makes it a safe bet to increase the number of iterations if one is unsure whether convergence has been reached.
-Across all cases examined, we found convergence to be quite fast.
-Almost without exception, 100 iterations were more than enough, by a large margin.
+6.1 Choice of Hyperparameters
+mini-vec2vec is stable under different choices of parameters, often converging to similar stationary points.
+Moreover, unlike many machine learning algorithms, most hyperparameters have a clear interpretation and often have a monotone behavior.
+Here we share trends that we have observed empirically in some setups, but a rigorous experimental analysis has not been made.
+  • Number of concatenated representations: Increasing the number of runs in the anchor discovery step reduces the number of the failure modes.
+This is because noisy assignments are diluted, emphasizing the signal from true assignments.
+Note that even with very few runs, good alignments can be reached at random but with smaller probability.
+  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows, unlike some algorithms where increasing information might cause degradation.
+This makes it a safe bet to increase the number of iterations.
+Across all cases examined, though, we found convergence to be quite fast.
+Almost without exception, 100 iterations were enough, by a large margin.
   • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise.
 The larger the sample size is, the closer this approximation is to the data.
-We generally consider higher sampling rates to be better.
+As a rule of thumb, we generally considered higher sampling rates to be better (at the cost of slower performance), but we haven’t tested this rigorously.
 Other hyperparameters are less obvious.
 Number of clusters in the approximate matching step is likely good in moderation.
-While it gives us more landmark points for alignment, it can theoretically lead to instabilities, because the assignment of very granular centroids can be noisy.
-Increasing number of neighbors is often beneficial, to a point.
-It helps reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position.
-Note that this is in contrast to multilingual word embeddings, where essentially every point has a ground-truth translation, and it makes sense to rely on a single neighbor.
-Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance eventually.
-Finally, increasing the exponential smoothing parameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores.
-This is likely due to overfitting.
-The smoothing parameter is implicitly responsible for the maximum distance of the matrix from being orthogonal.
-While this (alleged) overfitting phenomenon has been observed by us empirically, we did not explore it rigorously, and it might not be a common case.
+While increasing it will give us more landmark points for alignment, it can theoretically lead to instabilities because the assignment of very granular centroids can be noisy.
+Increasing number of neighbors is often beneficial, but it eventually leads to diminishing returns, and might even deteriorate performance when neighborhoods become very large.
+At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position.
+Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation, and it makes sense to rely on a single neighbor.
+Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance.
+Finally, increasing $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores.
+$\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal.
 Additionally, increasing $\alpha$ makes convergence slightly slower.
-Unlike many deep learning setups – GANs in particular – the range of hyperparameters that lead to strong performance is generously wide.
-Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter in order to improve effectiveness, if needed.
-Failure Modes
-We find that despite the many experiments we ran, including ones with different configurations, failure modes were extremely rare.
-In fact, we didn’t experience any failure modes except with the gte model.
-Moreover, all failure modes were around the gte-gtr pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in gtr $\rightarrow$ gte – they are just rarer.
+In summary, unlike many deep learning setups, GANs in particular, the range of hyperparameters that lead to strong performance is generously wide.
+Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter to improve effectiveness, if needed.
+6.2 Failure Modes
+We find that despite the many experiments we ran, ranging diverse hyperparameter configurations, failure modes were extremely rare.
+In fact, we didn’t experience any except on the $gte\leftrightarrow gtr$ pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in the gtr $\rightarrow$ gte direction, they are just rarer.
 Crucially, failure modes are not catastrophic.
-The failure modes are characterized by worse top-1 accuracy and average rank (and slightly lower cosine similarity throughout the refinement steps.
-Often also noticeable in the initial step, but not always).
-But the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
+Failure modes are characterized by worse top-1 accuracy and average rank – but the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
+6.3 Alternative Approaches
+Through anecdotal experiments, we have found that it might be possible to sidestep the use of relative representations in the approximate matching step.
+Instead, we can perform Procrustes Analysis directly on the concatenated aligned centroids.
+It seems to do slightly worse in the approximate matching step, but it converges normally after.
+We emphasize, as noted above, that we haven’t tested this thoroughly to be certain it generalizes.
+For future reference, we call this variant mini-vec2vec without relative representations”, but catchier names can be used if meaning is clear.
 7 Related Work
@@
 Representation alignment is well-studied in the supervised setup.
-Lenc and Vedaldi (2015) and Bansal et al.
-(2021) demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach.
-Olah (2015) studied the equivalence of neural networks in relative terms.
-Moschella et al.
-(2023) have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching.
-Similarly, Dar et al.
-(2023) have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors.
-Huh et al.
-(2024) have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
-The task of unsupervised alignment of embeddings has a long history in word embeddings (Conneau et al., 2018; Chen and Cardie, 2018; Grave et al., 2018) and machine translation (Lample et al., 2018; Artetxe et al., 2018).
-Most of the works rely on training a GAN.
-They often require a small seed of matched pairs in order to work.
-Hoshen and Wolf (2018) avoid the adversarial setup and use a three-step algorithm similar to ours.
-Their approach, specifically their approximate matching process, which relied on PCA, was insufficient for our purposes.
+lenc2014stitching and bansal2021revisiting have demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach.
+colah2015visualizing studied the equivalence of neural networks in relative terms.
+moschella2022relative have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching.
+Similarly, dar2023analyzing have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors.
+huh2024platonic have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
+The task of unsupervised alignment of embeddings also has a long history in word embeddings (conneau2017word; chen-cardie-2018-unsupervised; grave2018unsupervisedalignmentembeddingswasserstein) and machine translation (lample2018unsupervisedmachinetranslationusing; artetxe2018unsupervisedneuralmachinetranslation).
+Most of the methods rely on training a GAN, and they often require a small seed of matched pairs in order to work.
+hoshen-wolf-2018-non avoid the adversarial setup and use a three-step algorithm similar to ours.
+Their approach, especially the approximate matching step, which relied on PCA, was insufficient for our purposes.
 8 Conclusion
 We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability.
-Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, provided that structural correspondences can be estimated accurately.
-The success of our approach provides strong empirical evidence for the universal geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed.
+Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways.
+The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed.
 By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
-The efficiency and accessibility of our method could democratize research on embedding alignment and cross-modal representation learning.
+The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning.
 By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments.
 From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces.
-If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al.
-(2025).
+If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in jha2025vec2vec.
 Acknowledgements

### v4

@@
 1 Introduction
-Sentence embeddings capture rich semantic structure in high-dimensional vector spaces.
-They encode complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval.
-However, different architectures, training procedures, and even different models with the same architecture produce embedding spaces that cannot be compared directly, because they use a different coordinate system, a different “internal language”.
-The problem of aligning embedding spaces without parallel supervision – that is, without pairs of matching encodings of the same data – has emerged as a fundamental challenge with far-reaching implications.
-Successful unsupervised alignment has multiple implications across many domains: multilingual natural language processing (conneau2017word; zhang-etal-2017-adversarial), cross-modal alignment between text and image representations (maniparambil2024; schnaus2025), and privacy concerns (song2020informationleakage; text2vec).
-The recent vec2vec method (jha2025vec2vec) has demonstrated that unsupervised alignment is indeed possible using CycleGAN (zhu2017cyclegan).
-Their approach is theoretically grounded in a principle called the Universal Geometry or the Platonic Representation Hypothesis (huh2024platonic), which posits that well-trained representations, regardless of their specific architectures or training procedures, tend to converge toward geometrically similar spaces, even if their coordinate systems differ.
-This means that there is a way to rotate the coordinate systems so they “understand” each other – and vec2vec does precisely that; it identifies a rotation that makes the coordinate systems aligned.
-While vec2vec has proven effective, adversarial training frameworks like CycleGAN bring inherent challenges that limit their practical applicability.
-Training them is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (saxena2023gan_survey; goodfellow2017nips2016).
-Moreover, they typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
-Our point of departure is still Universal Geometry, but we take a different approach.
-We are motivated by the remarkable success of supervised alignment approaches, particularly neural network stitching methods (lenc2014stitching; bansal2021revisiting), which have demonstrated that simple affine transformations often suffice to align representations.
-But unfortunately, stitching assumes access to pairs of matching representations.
-Motivated by the Platonic Representation Hypothesis, we hypothesize that we can nonetheless uncover this linear alignment without even needing adversarial training.
-Specifically, we hypothesize that the similarity in the geometric structure – the relative arrangement of semantic clusters, the preservation of neighborhood structures, and the overall geometric organization of concepts – should provide sufficient signal for alignment without parallel supervision and without need for adversarial methods.
-To this end, we utilize the Relative Representations framework of moschella2022relative, which allows us to operate in a universal space agnostic of arbitrary coordinate systems.
-We then use it as a stepping stone to produce a noisy alignment that we refine with further steps.
-[^1] In Section 6, we discuss a variant that sidesteps the use of relative representations, but we tested it far less.
-mini-vec2vec is a simple and robust pipeline with competitive alignment performance offering several key advantages over adversarial approaches: computational efficiency (very fast and running on commodity CPU hardware), training stability (good performance and robustness to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (effective with substantially fewer training examples).
+Text embeddings, also called sentence embeddings, represent chunks of text in a high-dimensional vector space.
+They capture complex linguistic relationships and enable downstream tasks, such as semantic similarity and information retrieval.
+Naturally, different embedding models induce incompatible embedding spaces that cannot be compared directly, due to differences in architecture or initialization.
+The output of one model looks like gibberish to another.
+This can be easily solved with supervised training – the translation between embedding models can be learned from embedding pairs representing the same sentences.
+However, the unsupervised problem, where we don’t have access to parallel embeddings of the same sentences, is much harder.
+The unsupervised alignment of word embeddings has been studied extensively for a long time, often for the practical goal of aligning monolingual word embeddings from different languages.
+However, text embeddings have been mostly ignored.
+With the increasing popularity of databases of text embeddings, leakage of sensitive vector data becomes a real possibility.
+It has been shown that embedding inversion methods can extract the texts from the vectors alone (Song and Raghunathan, 2020; Morris et al., 2023), but they require access to the model.
+Unsupervised alignment can be used to extend the attack to unknown proprietary models.
+The attacker just translates the embeddings to a known space and applies inversion methods on the translated embeddings – extracting the text from the data (Jha et al., 2025).
+Recently, Jha et al.
+(2025) have proposed vec2vec, a method based on CycleGAN (Zhu et al., 2017) for fully unsupervised text embedding alignment.
+As the logic behind their method, they cite a principle sometimes called Universal Geometry or the Platonic Representation Hypothesis (Huh et al., 2024), which posits that well-trained representations, regardless of architectures or initialization, tend to converge toward geometrically similar spaces.
+This means that there is a way to rotate one embedding space to make it compatible with the other, and vice versa.
+While vec2vec is successful in finding a good alignment, the use of adversarial training is computationally intensive, requires careful hyperparameter tuning, and is subject to training instabilities like mode collapse, oscillatory behavior, or failure to converge (Saxena and Cao, 2023; Goodfellow, 2017).
+Moreover, adversarial methods typically demand substantial computational resources, including GPU acceleration and large amounts of training data – factors that may prohibit their utilization by most researchers and users.
+We hypothesize that the transformation between the embedding spaces is of a simple form (linear, specifically), and that it can be learned without adversarial methods.
+Indeed, in word embedding alignment much of the work has focused on this setup as well (Artetxe et al., 2018; Hoshen and Wolf, 2018; Grave et al., 2019, inter alia).
+To achieve a similar goal in sentence embeddings, mini-vec2vec seeks to identify signals of similarities in the geometric structure of sentence embedding spaces.
+We propose a very simple pipeline (see Figure 1 for a schematic overview), which consists of three steps:
+  • Approximate Matching: We match, in an unsupervised way, a small set of artificial pairs.
+  • Initial Transformation: We find a linear map between the pairs and apply it to the full data.
+  • Iterative Refinement: We then iteratively apply the operator, match each embedding to the closest embeddings in the other space, and learn a new operator on top of this matching.
+This pipeline resembles some methods from unsupervised word embedding alignment – but text embeddings are somewhat different from word embeddings, so they require other tricks to make it work.
+Perhaps the most significant difference is the Approximate Matching method we rely on.
+At a high level, we need to find a new way to identify stable central landmarks and match them between sentence embedding spaces.
+Methods in word alignment implicitly rely on word vocabularies, which have special features (e.g., common words between languages, word frequencies, distributional properties of words, etc.).
+On the other hand, sentence embeddings can be seen as sparse samples from an “infinite” vocabulary, rendering them closer to, for instance, image embeddings.
+After we have found an approximate match, the differences between word and sentence embeddings blur, and similar methods can be applied.
+We have developed the pipeline unaware of existing literature, so it is possible that our design decisions presented here are suboptimal or less common.
+Having said that, even in the Initial Alignment and Iterative Refinement steps, the design caters to specific characteristics uncommon in word embeddings, or at least less salient.
+For instance, when matching embeddings to their nearest neighbors, we average multiple nearest neighbors because there is no one-to-one matching between data points.
+This is often not the case in word alignment, since most languages, even if they are very distant, rely on closed vocabularies that significantly overlap.
+We find mini-vec2vec offers several key advantages: computational efficiency (very fast and runs on a commodity CPU), training stability (good performance and robust to hyperparameters), interpretability (each step has a clear geometric interpretation), and sample efficiency (60k vs 2 million samples).
 Through comprehensive experiments, we demonstrate that this simple approach not only matches but often exceeds the performance of the more complex adversarial method while requiring orders of magnitude less computational resources.
+Figure 1: A simplified illustration of the mini-vec2vec pipeline.
 2 Preliminaries
 The hypothesis that neural networks converge to geometrically similar representations has been explored across multiple domains and scales.
-Early work by colah2015visualizing has revealed striking structural similarities in representations of different convolutional networks, [^2] Olah briefly refers to even earlier ideas (e.g., a concept called “meta-tSNE” by Hinton), but we failed to locate them.
-suggesting that the geometric organization of learned features might be more universal than previously assumed.
-The Platonic Representation Hypothesis, formally articulated by huh2024platonic, represents the most comprehensive theoretical framework for understanding Universal Geometry.
+Early work by Olah (2015) has revealed striking structural similarities in representations of different convolutional networks, suggesting that the geometric organization of learned features might be more universal than previously assumed.
+The Platonic Representation Hypothesis, formally articulated by Huh et al.
+(2024), represents the most comprehensive theoretical framework for understanding Universal Geometry.
 The hypothesis posits that well-trained representations converge toward a shared geometric structure that reflects the underlying structure of the data domain.
-Having the same geometry implies that the distance between objects should be independent of the models themselves.
-For example, if one wants to compute the semantic similarity of “I love your dog” and “I love my cat” using cosine similarity between their embeddings, the answer should be roughly the same regardless of the underlying embedding model.
-Intuitively, this empirical observation stands to reason, as semantic similarity originates from the semantic meaning that comes from the data itself.
-Below, we discuss the framework of moschella2022relative that extends this observation into a powerful tool.
+This means that distances (and similarities) between objects should be roughly independent of the embedding model.
+For example, to compute the semantic similarity between “I love your dog” and “I love my cat”, we would use cosine similarity between their embeddings.
+According to Universal Geometry, the answer should (theoretically) have little dependence on the model used.
+This makes sense intuitively, as semantic similarity originates from semantic meaning.
+It is a fundamentally linguistic concept learned from the data itself.
 Relative Representations
-moschella2022relative have operationalized the observations of colah2015visualizing and proposed to use relative representations as a mathematically principled framework for comparing latent spaces.
-While absolute coordinates vary arbitrarily across models, we know that relative information, such as similarity between two objects, is (approximately) agnostic to the choice of embedding model.
-They propose to use a shared set of anchor points that other data points will be compared to as a way to ground the position of data points in a model-agnostic way.
-Denote the space of data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality) and the embedding function by $E:\mathcal{X}\rightarrow\mathbb{R}^{d}$.
-Given a set of anchor data points $\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under $E$ is:
-$\mathbf{r}_{E}(\mathbf{z})=[f(E(\mathbf{z)},E(\mathbf{a}_{1}));f(E(\mathbf{z}),E(\mathbf{a}_{2}));\ldots;f(E(\mathbf{z}),E(\mathbf{a}_{k}))],$
-where $f(\cdot,\cdot)$ is a similarity function (typically cosine similarity) and ; represents concatenation.
-The power of relative representations is that, as shown by colah2015visualizing and moschella2022relative, for two encoders $E_{1},E_{2}$, the relative representations of the same data point is roughly the same, i.e., $\mathbf{r}_{E_{1}}(\mathbf{z})\approx\mathbf{r}_{E_{2}}(\mathbf{z})$.
-In other words, the function $\mathbf{r}_{E}:\mathcal{X}\rightarrow\mathbb{R}^{D}$ is itself an embedding that maps a data point into $k$-dimensional space in a way that is approximately independent of $E$ (for a sufficiently well-trained $E$) and only depends on the choice of anchor points.
-3 Problem Formulation
-We follow the setup laid out in vec2vec.
-We assume access to two pools of text embeddings, each from a different encoder, and our task is to align their embedding spaces without paired examples.
-Similar to vec2vec, we learn a parameterized mapping (neural network or linear transformation) rather than a one-to-one matching between data points.
-This approach offers several advantages:
-  • Generalization: New embeddings can be translated without re-applying the algorithm
-  • Robustness: Works when one-to-one point matching doesn’t exist, requiring only similar data distributions
-  • Out-of-distribution capability: The parameterized nature can (potentially) enable generalization beyond the training distribution
-The Universal Geometry hypothesis indicates that such a task can be possible: “good enough” models learn equivalent geometries – relative similarities between data points remain approximately the same regardless of the embedding model.
-Let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings, where $\mathbf{x}^{A}_{i},\mathbf{x}^{B}_{j}\in\mathbb{R}^{d}$.
-The sentences come from the same underlying distribution $\mathcal{D}$, but – unlike most other works – there is no overlap between the sentences themselves – only the distribution is shared.
-We organize these embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector.
-The fundamental challenge is that while both spaces may contain similar geometric structures, they are related by an unknown transformation.
+Moschella et al.
+(2022) have operationalized the observations of Olah (2015) and proposed to use relative representations as a mathematically principled framework for comparing latent spaces.
+While absolute coordinates vary arbitrarily across models, we know now that relative information, like similarity between two objects, is (approximately) agnostic to the choice of embedding model.
+They propose to choose an agreed-upon set of data points, called anchors, that other data points will be compared to, and use them like landmarks on a map.
+In order to specify the location of a point, we specify how close it is to every landmark.
+If we have enough anchors and they are sufficiently diverse, we can “geo-locate” the point by its relative position alone.
+Since it requires a small set of paired anchors, the method cannot operate as-is in a fully unsupervised manner.
+Denote the set of possible data points by $\mathcal{Z}$ (e.g., sentences, images, or any other modality; note these are abstract objects and not their vector representations) and the embedding function by $E:\mathcal{Z}\rightarrow\mathbb{R}^{d}$.
+Given a set of anchor data points $A=\{\mathbf{a}_{i}\}_{i=1}^{k}\subseteq\mathcal{Z}$, the relative representation of $\mathbf{z}\in\mathcal{Z}$ under anchors $A$ and embedding $E$ is the vector:
+$\mathbf{r}_{A,E}(\mathbf{z})=[\,\operatorname{sim}(E(\mathbf{z)},E(\mathbf{a}_{1})),\,\operatorname{sim}(E(\mathbf{z}),E(\mathbf{a}_{2})),\,\ldots,\,\operatorname{sim}(E(\mathbf{z}),E(\mathbf{a}_{k}))\,],$
+where $\operatorname{sim}(\cdot,\cdot)$ is a similarity function (typically cosine similarity).
+For two embedding models satisfying the Universal Geometry assumption, relative representations of the same data point are roughly the same, i.e., $\mathbf{r}_{A,E_{1}}(\mathbf{z})\approx\mathbf{r}_{A,E_{2}}(\mathbf{z})$.
+3 Methodology
+Problem Formulation
+Following the setup laid out in vec2vec, let $\mathcal{X}_{A}=\{\mathbf{x}^{A}_{i}\}_{i=1}^{n_{A}}$ and $\mathcal{X}_{B}=\{\mathbf{x}^{B}_{j}\}_{j=1}^{n_{B}}$ denote two sets of $d$-dimensional embeddings.
+The sentences that created these embeddings are drawn from a shared underlying distribution $\mathcal{D}$, but all sentences are different and there is no overlap between the two datasets.
+We organize the embeddings into matrices $\mathbf{X}_{A}\in\mathbb{R}^{n_{A}\times d}$ and $\mathbf{X}_{B}\in\mathbb{R}^{n_{B}\times d}$, where each row corresponds to a single embedding vector.
 Our goal is to learn a transformation function $f:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that $f(E_{A}(\mathbf{z}))\approx E_{B}(\mathbf{z})$ where $\mathbf{z}$ represents a data point $\mathbf{z}\in\mathcal{Z}$.
-More generally, the setup described in jha2025vec2vec allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$.
-We will not take advantage of this flexibility in this paper.
-[^3] It also complicates the analysis of the results.
-For example, $f_{A}=f_{B}=\text{const}$ would be a trivial solution if we don’t impose additional constraints.
-Similarity then needs to be assessed more carefully.
-4 Methodology
-4.1 Motivation
-In this work, we restrict our attention to linear transformations, i.e., $f(\mathbf{x})=\mathbf{W}\mathbf{x}$ where $\mathbf{W}\in\mathbb{R}^{d\times d}$.
-Moreover, we will constrain our search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
-This is motivated by the Universal Geometry hypothesis – we expect cosine similarities to be preserved across embedding spaces.
-Later on, we will relax this assumption and ultimately the matrix will be an exponentially-smoothed average of orthogonal matrices, enforcing orthogonality softly.
-Our linear approach, mini-vec2vec, addresses key limitations of the current approach: GANs are notoriously unstable and hard to train, demand extensive computational resources and training time.
-They often require tedious hyperparameter tuning and sometimes multiple re-runs to converge to good solutions.
-On the other hand, we know that in supervised cases (model stitching), there exists an affine mapping between the latent spaces, suggesting the linear case is realizable – the question is how to find it efficiently.
-mini-vec2vec achieves that, and it has several appealing properties: it requires significantly less data than vec2vec (60k vs.
-1 million samples), runs on a CPU rather than requiring extensive GPU resources, and provides more stable training dynamics.
-Our method resembles the approach described in hoshen-wolf-2018-non for word embeddings, proceeding in three main stages: approximate matching, transformation fitting, and iterative refinement.
-This paradigm is very powerful, and we adopt a similar approach.
-Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because there is no small exhaustive set of objects (words) that need to be matched.
-Moreover, the very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small.
-Therefore, word embedding alignment is rarely considered in the non-overlapping datasets setup (even in multilingual matching, words mostly have a ground-truth translation).
-We find that applying the algorithm from hoshen-wolf-2018-non naively to our use case doesn’t work.
-The approximate matching stage, in particular, requires more delicate treatment in our case.
-The iterative refinement step is also improved by a slightly more subtle approach.
+More generally, the setup described in Jha et al.
+(2025) allows for learning a pair of functions $f_{A},f_{B}$ sending embeddings to a third space, such that $f_{A}(E_{A}(\mathbf{z}))\approx f_{B}(E_{B}(\mathbf{z}))$, but we will not take advantage of this flexibility here.
+Motivation
+In this work, we restrict our attention to linear transformations.
+Further, we constrain the search space to orthogonal transformations, i.e., matrices $\mathbf{W}$ that satisfy $\mathbf{W}\mathbf{W}^{T}=\mathbf{W}^{T}\mathbf{W}=\mathbf{I}$.
+This is motivated by the Universal Geometry hypothesis – we expect distances and similarities to be preserved by the transformation.
+In practice, we later relax this assumption and use an exponentially-smoothed average of orthogonal matrices to enforce orthogonality softly.
+mini-vec2vec resembles many approaches in unsupervised word embedding alignment.
+It proceeds in three main stages: approximate matching, transformation fitting, and iterative refinement.
+Unfortunately, text embeddings are somewhat more complicated than word embeddings, partly because they are non-overlapping – there is no small exhaustive set (a vocabulary) of objects that need to be matched.
+The very concept of non-overlapping datasets is somewhat contrived for word embeddings, precisely because the set of possible objects is closed and relatively small.
+Even in multilingual matching, words mostly have a ground-truth one-to-one match.
+Beyond the overlap between the datasets, most works (Artetxe et al., 2016, 2017, 2018; Hoshen and Wolf, 2018) rely on the statistics of words or other word-specific features (for finding a seed dictionary).
+Optimal transport-based methods (Alvarez-Melis and Jaakkola, 2018; Grave et al., 2019) rely heavily on matching points between the sets (often one-to-one, but this can be alleviated), and they become computationally intensive.
+Theoretically speaking, in a non-overlapping dataset, OT could find sentences that are (only) similar and match them to one another, and then a parametric mapping could be learned on top of them (a methodology similar to Grave et al.
+(2019), only applied to non-overlapping data).
+In practice, when Jha et al.
+(2025) have compared their method to multiple OT methods, in an oracle setup, they didn’t work better than the most naive baseline, indicating that even in the oracle setup, and even as a first step for learning a mapping, they are insufficient.
+We will now proceed to describe our algorithm in detail (pseudo-code is provided in Listing LABEL:lst:full_algo).
 def mini_vec2vec(X_A, X_B):
@@
 C_B = k_means(X_B, num_clusters)
-S_A = similarity_function(C_A, C_A)
-S_B = similarity_function(C_B, C_B)
+S_A = pairwise_similarities(C_A, C_A)
+S_B = pairwise_similarities(C_B, C_B)
 pi = solve_QAP(S_A, S_B)
 C_B_permuted = permute(C_B, pi)
-R_A_list.append(similarity_function(X_A, C_A))
-R_B_list.append(similarity_function(X_B, C_B_permuted))
+R_A_list.append(pairwise_similarities(X_A, C_A))
+R_B_list.append(pairwise_similarities(X_B, C_B_permuted))
 R_A, R_B = concatenate(R_A_list), concatenate(R_B_list)
@@
 Listing 1: Pseudo-code for mini-vec2vec (refinement algorithms deferred to Listing LABEL:lst:refine)
-We will now describe our algorithm in detail. Pseudo-code is provided in Listing LABEL:lst:full_algo.
 3.1 Preprocessing
@@
 $\hat{\mathbf{X}}_{A}[i,:]$ $=\frac{\tilde{\mathbf{X}}_{A}[i,:]}{\|\tilde{\mathbf{X}}_{A}[i,:]\|_{2}},\quad\text{similarly for space }B$
-where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ denotes the L2 norm.
+where $\boldsymbol{\mu}_{A}$ is the mean of space $A$, $\mathbf{1}_{n_{A}}$ is the vector of ones of length $n_{A}$, and $\|\cdot\|_{2}$ is the L2 norm.
 3.2 Approximate Matching
-To establish structural correspondences, we employ an anchor discovery procedure based on multiple runs of clustering and matching.
-We will first describe a single run, and then explain how we robustify it by ensembling multiple runs:
+To find an approximate matching, a small seed of pairs of (approximate) matches, we need to find stable landmarks that are likely to exist in both spaces, despite the non-overlapping data.
+The natural candidates are clusters.
+We conjecture that clusters represent recurring themes between the datasets and, therefore, should be consistent between the datasets.
+But even if we find the same set of clusters in both spaces exactly, we still need to match each cluster (centroid) to its corresponding cluster (centroid) in the other space.
+To match the clusters, we rely on pairwise similarities, conjecturing they encode enough information to identify the matching uniquely (this intuition is shared with Gromov-Wasserstein (Mémoli, 2011)).
+We match the clusters based on similarity matrices, with Quadratic Assignment Problem (QAP) solvers.
+These assumptions are not trivial, but the method works relatively well as-is, though often requiring a few attempts due to the stochasticity and sensitivity of the clustering and matching algorithms.
+To mitigate this, we run cluster-and-match multiple times and ensemble the results.
+We will now describe a single run, and then explain how we robustify it by ensembling.
 Centroid Matching
-The reason we choose to align cluster centroids is that they roughly represent the central landmarks in the dataset.
-Since the data comes from a shared distribution $\mathcal{D}$, we expect clusters to be similar overall, and all that remains to do is to find the alignment between them.
-Of course, clusters are not very consistent even with exactly the same data, so the approach needs some tolerance to errors (hence the robustification step below).
-We perform k-means clustering on both embedding spaces independently to obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,k$.
-We then compute similarity matrices between the centroids independently in each embedding space:
+We perform k-means clustering in each embedding space independently and obtain cluster centroids $\{\mathbf{c}^{A}_{j}\}$ and $\{\mathbf{c}^{B}_{j}\}$ where $j=1,\ldots,C$.
+We then compute pairwise similarities in each space:
 $S^{A}_{ij}$ $=\cos(\mathbf{c}^{A}_{i},\mathbf{c}^{A}_{j}),\quad S^{B}_{ij}=\cos(\mathbf{c}^{B}_{i},\mathbf{c}^{B}_{j})$
 where $\mathbf{c}^{A}_{i}$ and $\mathbf{c}^{B}_{i}$ are the $i$-th cluster centroids in spaces $A$ and $B$ respectively.
-We then find the optimal matching between cluster centroids from different embedding spaces by solving the Quadratic Assignment Problem (QAP):
-$\pi^{*}=\arg\max_{\pi\in\Pi_{k}}\;\sum_{i=1}^{k}\sum_{j=1}^{k}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
-where $\pi:\{1,2,...,k\}\rightarrow\{1,2,...,k\}$ is a permutation from the set $\Pi_{k}$ of all permutations on $k$ elements.
+We then find the optimal matching between the cluster centroids by solving the Quadratic Assignment Problem (QAP), which finds a permutation that aligns the similarity matrices optimally:
+$\pi^{*}=\arg\max_{\pi\in\Pi_{C}}\;\sum_{i=1}^{C}\sum_{j=1}^{C}S^{A}_{ij}\,S^{B}_{\pi(i)\,\pi(j)}$
+where $\pi:\{1,2,...,C\}\rightarrow\{1,2,...,C\}$ is a permutation from the set $\Pi_{C}$ of all permutations on $C$ elements.
 Due to the small number of cluster centroids, this takes only a few seconds.
-Relative Representations
+Conceptually, we can think of QAP in terms of cosine similarity, where we want to find the permutation that will maximize the cosine similarity between the (flattened) similarity matrices.
+Option 1: Relative Representations
 Once we have identified an alignment between $A\text{-}$ and $B\text{-centroids}$, we use the aligned centroids as anchors.
@@
 For embedding $\mathbf{x}^{A}_{i}$, we construct:
-$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{k})]$
+$\mathbf{r}^{A}_{i}=[\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{1}),\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{2}),\ldots,\cos(\mathbf{x}^{A}_{i},\mathbf{c}^{A}_{C})]$
 Similarly, for $\mathbf{x}^{B}_{i}$, we compute (note that we need to re-arrange the indices with $\pi$):
 […]
+Option 2: Alternative Approach
+Curiously, we have found that it is possible to sidestep relative representations. Instead, it is possible to use the aligned centroids directly.
+$\mathbf{r}^{A}_{i}=[\mathbf{c}^{A}_{1},\mathbf{c}^{A}_{2},\ldots,\mathbf{c}^{A}_{C}];\qquad\mathbf{r}^{B}_{i}=[\mathbf{c}^{B}_{\pi(1)},\mathbf{c}^{B}_{\pi(2)},\ldots,\mathbf{c}^{B}_{\pi(C)}]$
+While the relative representations method learns from information derived from the entire data, the centroid-only approach requires just the centroids themselves.
+Both methods are efficient, and this simplified method is only slightly more efficient.
+It seems to do worse after the Initial Transformation step (Section 3.3), but it converges normally in the Iterative Refinement steps (Section 3.4).
+We emphasize that we haven’t tested it as thoroughly as the relative representation approach.
+For future reference, we call this variant mini-vec2vec centroid-only”.
+Note that the centroid-only method is also much closer in spirit to many of the alignment methods discussed above that find a seed dictionary matching (the centroids here) and build on that.
+In this paper, we will not focus on this variant and discuss it here only for completeness.
 Robustification via Ensembling
 We improve robustness by repeating the process described above and concatenating relative representations from multiple runs:
-$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sk},\qquad\text{similarly for space }B$
-This concatenation creates richer structural signatures that are more robust to noise that comes from bad alignments.
-4.4 Mapping Estimation
+$\mathbf{r}^{A}_{i}=[\mathbf{r}^{A}_{i,1};\mathbf{r}^{A}_{i,2};\ldots;\mathbf{r}^{A}_{i,s}]\in\mathbb{R}^{sC},\qquad\text{similarly for space }B$
+This concatenation creates richer structural signatures that are more robust to the noise that comes from bad alignments.
+3.3 Initial Transformation
 We construct pseudo-parallel pairs by sending each element in space $A$ to the average of its $k$ nearest neighbors from space $B$ (based on similarity in relative space):
@@
 Refine-1: Matching-based Refinement
-This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.4), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space.
-The method is cheap and it runs for many iterations (50–100), using subsampling to boost computational efficiency further.
+This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Initial Transformation stage (Section 3.3), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space.
+This technique is also present in Hoshen and Wolf (2018), and it is often called Iterative Closest Point (ICP; Besl and McKay, 1992).
+A minor difference is that we average multiple points rather than relying on a single neighbor.
+As mentioned above, this is due to our no-overlap setup.
+Since ICP is cheap, we run it for many iterations (50–100), using subsampling to boost computational efficiency further.
 We use exponential smoothing for transformation updates:
@@
 Refine-2: Clustering-based Refinement
-This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids.
+This strategy aims to improve the large-scale matching between the spaces.
 We begin by clustering the embeddings in space $A$.
-We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization.
-The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors.
-The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids.
-This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$.
-Similar to the first strategy, we use exponential smoothing.
+We then apply $\mathbf{W}$ to the cluster centroids and cluster the $B$ embeddings, where the clustering algorithm is initialized with the transformed $A$ centroids as the initial centroids.
+The transformed clusters should be close to a set of clusters in the new space.
+We therefore expect the clustering algorithm to make only minor adjustments, correcting biases in the transformation and moving the centroids to their “right positions” in space $B$.
+We expect the pseudo-parallel pairs based on cluster assignments to be more robust than individual nearest neighbors used in Refine-1.
+Similar to the first refinement strategy, we use exponential smoothing.
 Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1.
-Two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
+Two or more lead to a slight deterioration of the alignment.
+We conjecture that the ICP procedure might have introduced systematic bias that requires correction, but after this correction, the clustering-based refinement does not contribute, and only dilutes the signal.
 def refine_1(X_A, X_B, W):
@@
 4 Experiments
-We replicate the main experiment from jha2025vec2vec.
-We run our algorithm for all pairs of encoders used in jha2025vec2vec.
-Details about the encoders are provided in Table 2.
-Experiments use k-means clustering from scikit-learn (scikit-learn) and QAP is solved with the scipy (2020SciPy-NMeth) implementation of 2-OPT.
+We replicate the main experiment from Jha et al.
+(2025).
+We run our algorithm for all pairs of encoders used in Jha et al.
+(2025).
+Details about the encoders are provided in Table 2 in the Appendix.
+Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011) and QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT.
 We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score.
 This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
-We take a random subset of the Natural Questions dataset (natural_questions) of size $60,000$ and compute sentence embeddings under all possible encoders.
-We leave $8192$ sentences for evaluation.
-For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences.
-All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate consistency.
-We compare our results to the numbers reported in jha2025vec2vec.
+We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$, compute sentence embeddings with all encoders, and leave $8192$ sentences for evaluation.
+We split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences.
+All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate the method’s stability.
+We compare our results to the numbers reported in Jha et al.
+(2025).
 Metrics
@@
 We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range.
-We make an economical choice of hyperparameter values due to the large number of runs, but we find the choice to be sufficiently robust still.
+For the paper, we made an economical choice of hyperparameter values to shave time for the large number of runs, but we found the choice to be sufficiently robust even still.
 We use $s=30$ runs of relative representation alignment with $c=20$ clusters.
-In Mapping Estimation, we use $k=50$ neighbors.
+In the Initial Transformation step, we use $k=50$ neighbors.
 We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters.
@@
 We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
+A detailed discussion of observations about hyperparameter choice can be found in Appendix B.
 vec2vec mini-vec2vec
@@
 Results
-In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table.
+From the results in Table 1, we observe that our approach matches or exceeds vec2vec across (almost) the entire table.
 The improvement is most noticeable in the rank metric.
@@
 Analysis
-The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that it is achieved for a fraction of the cost.
+The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that this is achieved for a fraction of the cost.
 One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware.
 Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance.
-Analyzing the convergence of our algorithm, we observe that despite variation in the approximate matching step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations.
+Analyzing the convergence of our algorithm, we observe that despite variation in the Initial Transformation step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different initializations.
 This reproducible convergence behavior suggests the optimization landscape is well-structured.
 The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective.
-However, applying Refine-2 successfully improves upon this converged solution.
-The two methods appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
+Refine-2 successfully improves upon this converged solution, which might indicate that the converged solution has some systematic bias.
+5 Discussion
+Failure Modes
+We find that despite the many experiments we ran, ranging diverse hyperparameter configurations, failure modes were extremely rare.
+In fact, we didn’t experience any except on the $gte\leftrightarrow gtr$ pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in the gtr $\rightarrow$ gte direction, they are just rarer.
+Crucially, failure modes are not catastrophic.
+Failure modes are characterized by worse top-1 accuracy and average rank – but the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
+Main Takeaways
+We would like to emphasize the subtle points that make the method work.
+While word alignment methods weren’t suitable in this setup, they are likely not completely irrelevant.
+The main drivers of success are stable landmarks and tractable matching.
+We chose to focus on clusters as these landmarks.
+However, it is not the only option.
+A preliminary study (Yamagiwa et al., 2023) on word and image embeddings has found that Independent Component Analysis (ICA) can identify shared interpretable directions between embedding spaces.
+We don’t preclude the possibility that they could have been used instead.
+The important part is to find anchors that are consistent despite zero overlap.
+This gives us a match between sets, and leaves us with the task of aligning the individual anchors.
+The fact that the number of anchors is relatively small makes the process very fast, but also makes the problem simpler.
+It is easier to find near-optimal solutions when the number of samples is not large, while the problem quickly gets out of hand due to its combinatorial nature.
+We used QAP for matching, which is closely related to Optimal Transport.
+The small scale does more than speed up the process, but likely also improves the performance of the algorithm.
+We believe this might have worked with Gromov-Wasserstein and other OT algorithms as well, as long as the same principles were kept.
+Last but not least, robustness to noise is another crucial driver of success.
+While Hoshen and Wolf (2018) re-initialize their pipeline (500 times, which is very expensive, though parallelizable) to fight noise and arbitrariness, we use many noisy assignments in one run, addressing the problem without multiple runs.
+Other tricks, not necessarily common in literature, such as exponentially-smoothed orthogonal matrices, cluster-based matching, and others, have empirically contributed to boosting performance significantly, but even without them performance was decent.
+6 Related Work
+Universality
+The study of universality has its roots in several works.
+Representation alignment is well-studied in the supervised setup.
+Lenc and Vedaldi (2015) and Bansal et al.
+(2021) have demonstrated that linear transformations are often sufficient for aligning neural network representations in a procedure they called stitching.
+Mikolov et al.
+(2013) have identified that word embeddings from different languages converge to similar geometries and used it to align them with a learned linear transformation from a seed dictionary.
+Olah (2015) visualized representations from different convolutional neural networks using relative representations.
+Moschella et al.
+(2022) have used these insights to transition between representation spaces through a universal space, utilizing it for zero-shot stitching.
+Concurrently, Dar et al.
+(2023) have demonstrated zero-shot stitching using (a method equivalent to) relative representations with token embedding anchors.
+Huh et al.
+(2024) have aggregated and unified previous work under the name Platonic Representation Hypothesis.
+Unsupervised Alignment
+The task of unsupervised alignment of embeddings, too, has a long history in word embeddings (Conneau et al., 2017; Chen and Cardie, 2018; Zhang et al., 2017) and machine translation (Lample et al., 2017).
+These early methods relied on adversarial training.
+Artetxe et al.
+have started a line of supervised and semi-supervised alignment methods relying on an iterative refinement process (Artetxe et al., 2016, 2017), which culminated in VecMap (Artetxe et al., 2018), an unsupervised method that relied on the distributional features of words.
+Another unsupervised work (Hoshen and Wolf, 2018) used a three-step algorithm not unlike our own, but it required hundreds of re-initializations and used word frequencies.
+The works of Alvarez-Melis and Jaakkola (2018) and Grave et al.
+(2019) have utilized variants of optimal transport to align word embeddings.
+They require the existence of a match between points, at least approximately, in order to work.
+Works on unsupervised alignment in other domains, for example, image or cross-modal image-text alignment, have mostly focused on semi-supervised setups and overlapping datasets (Schnaus et al., 2025; Maniparambil et al., 2024), which inspired our initial attempts.
+7 Conclusion
+We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability.
+Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways.
+The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed.
+By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
+The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning.
+By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments.
+From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces.
+If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al.
+(2025).
+Acknowledgements
+We thank John X. Morris for useful feedback. We thank Google for making CPU and GPU notebooks freely available.
+Appendix A Additional Assets
 Model Params Backbone Paper Year
-gtr 110M T5 gtr 2021
-e5 109M BERT e5 2022
-gte 109M BERT gte 2023
-stella 109M BERT stella 2023
-granite 278M RoBERTa granite 2024
+gtr 110M T5 Ni et al. 2021
+e5 109M BERT Wang et al. 2022
+gte 109M BERT Li et al. 2023
+stella 109M BERT Zhang et al. 2023
+granite 278M RoBERTa Awasthy et al. 2024
 Table 2: Embedding models used in our experiments
-6 Discussion
-6.1 Choice of Hyperparameters
+Appendix B Choice of Hyperparameters
 mini-vec2vec is stable under different choices of parameters, often converging to similar stationary points.
@@
 Across all cases examined, though, we found convergence to be quite fast.
-Almost without exception, 100 iterations were enough, by a large margin.
+Almost without exception, 100 iterations were enough by a large margin.
   • Subsampling rate: Subsampling allows us to speed up the algorithm quite substantially, but it can increase noise.
@@
 At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position.
-Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation, and it makes sense to rely on a single neighbor.
+Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation and it makes sense to rely on a single neighbor.
 Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance.
-Finally, increasing $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores.
+Finally, increasing the exponential smoothing hyperparameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores.
 $\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal.
@@
 Each hyperparameter is interpretable, and one can plausibly guess in which direction to modify the parameter to improve effectiveness, if needed.
-6.2 Failure Modes
-We find that despite the many experiments we ran, ranging diverse hyperparameter configurations, failure modes were extremely rare.
-In fact, we didn’t experience any except on the $gte\leftrightarrow gtr$ pair, and modifying the number of clusters to $c=30$ solved this, though experimentation with more runs has shown that failure modes still exist in the gtr $\rightarrow$ gte direction, they are just rarer.
-Crucially, failure modes are not catastrophic.
-Failure modes are characterized by worse top-1 accuracy and average rank – but the decrease is not catastrophic, leading to $\sim 90\%$ top-1 accuracy and 5105-10 average rank, much better than any naive baseline.
-6.3 Alternative Approaches
-Through anecdotal experiments, we have found that it might be possible to sidestep the use of relative representations in the approximate matching step.
-Instead, we can perform Procrustes Analysis directly on the concatenated aligned centroids.
-It seems to do slightly worse in the approximate matching step, but it converges normally after.
-We emphasize, as noted above, that we haven’t tested this thoroughly to be certain it generalizes.
-For future reference, we call this variant mini-vec2vec without relative representations”, but catchier names can be used if meaning is clear.
-7 Related Work
-The study of universality has its roots in several works.
-Representation alignment is well-studied in the supervised setup.
-lenc2014stitching and bansal2021revisiting have demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach.
-colah2015visualizing studied the equivalence of neural networks in relative terms.
-moschella2022relative have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching.
-Similarly, dar2023analyzing have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors.
-huh2024platonic have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
-The task of unsupervised alignment of embeddings also has a long history in word embeddings (conneau2017word; chen-cardie-2018-unsupervised; grave2018unsupervisedalignmentembeddingswasserstein) and machine translation (lample2018unsupervisedmachinetranslationusing; artetxe2018unsupervisedneuralmachinetranslation).
-Most of the methods rely on training a GAN, and they often require a small seed of matched pairs in order to work.
-hoshen-wolf-2018-non avoid the adversarial setup and use a three-step algorithm similar to ours.
-Their approach, especially the approximate matching step, which relied on PCA, was insufficient for our purposes.
-8 Conclusion
-We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability.
-Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways.
-The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed.
-By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
-The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning.
-By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments.
-From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces.
-If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in jha2025vec2vec.
-Acknowledgements
-We thank John X. Morris for useful feedback. We thank Google for making CPU and GPU notebooks freely available.